### v1

# Haoran Guan

Abstract
In this work, we focus on improving LU-CholeskyQR2 LUChol . Compared to other deterministic and randomized CholeskyQR-type algorithms, it does not require a sufficient condition on $\kappa_{2}(X)$ for the input tall-skinny matrix $X$, which ensures the algorithm’s safety in most real-world applications. However, the Cholesky factorization step may break down when the $L$-factor after the LU factorization of $X$ is ill-conditioned. To address this, we construct a new algorithm, LU-Householder CholeskyQR2 (LHC2), which uses HouseholderQR to generate the upper-triangular factor, thereby avoiding numerical breakdown. Moreover, we utilize the latest sketching techniques to develop randomized versions of LHC: SLHC and SSLHC. We provide a rounding error analysis for these new algorithms. Numerical experiments demonstrate that our three new algorithms have better applicability and can handle a wider range of matrices compared to LU-CholeskyQR2. With the sketching technique, our randomized algorithms, SLHC2 and SSLHC3, show significant acceleration over LHC2. Additionally, SSLHC3, which employs multi-sketching, is more efficient than SLHC2 and exhibits better numerical stability. It is also robust as a randomized algorithm.
Keywords:
QR factorization Rounding error analysis LU-CholeskyQR Numerical linear algebra
MSC:
65F30 15A23 65F25 65G50
1 Introduction
QR factorization is one of the most important components of numerical linear algebra and is widely used in both academia and industry (see 2018robust ; halko2011 ; 2010 ; 2020 ; 1989 for more details). Among the various algorithms for QR factorization, CholeskyQR has gained popularity in recent years due to its ability to balance efficiency and accuracy. CholeskyQR exclusively utilizes BLAS3 operations and requires only simple reductions in parallel environments, which is a significant advantage compared to other algorithms such as HouseholderQR, CGS (MGS), and TSQR ballard2011 ; 2011 ; MatrixC ; Higham ; Communication ; Numerical .
1.0.1 Deterministic CholeskyQR-type algorithms
CholeskyQR primarily focuses on the QR factorization of tall-skinny matrices. $X\in\mathbb{R}^{m\times n}$ and $m\geq n$, the Gram matrix $B$ is computed first, and the upper-triangular $R$-factor is obtained through Cholesky factorization. The orthogonal factor $Q\in\mathbb{R}^{m\times n}$ can then be derived using $Q=XR^{-1}$. CholeskyQR is outlined in Algorithm 1. Due to the requirement for Cholesky factorization in the algorithm, the input $X$ must satisfy $rank(X)=n$, which is a sufficient condition for all CholeskyQR-type algorithms.
Algorithm 1 $[Q,R]=\mbox{CholeskyQR}(X)$
1:  $B=X^{\top}X,$
2:  $R=\mbox{Cholesky}(B),$
3:  $Q=XR^{-1}.$
In practical applications, CholeskyQR is seldom used alone due to its lack of orthogonality. To enhance numerical stability in both orthogonality and residuals, CholeskyQR2 2014 ; error was developed by repeating CholeskyQR twice. This algorithm is outlined in Algorithm 2. To address the issue of ill-conditioned matrices, Shifted CholeskyQR (SCholeskyQR) Shifted introduces a shifted term $s$ during the Cholesky factorization step to avoid numerical breakdown, as detailed in Algorithm 3. Researchers have combined CholeskyQR2 with Shifted CholeskyQR to create Shifted CholeskyQR3 (SCholeskyQR3), which ensures orthogonality, as shown in Algorithm 4. In our recent work Columns , we define a new $g$-norm of $X$ and improve the original shifted term associated with $\|X\|_{2}$ by using a smaller term based on $\|X\|_{g}$. This modification expands the computational range of $\kappa_{2}(X)$ for Shifted CholeskyQR3. With the definition of the $g$-norm, we also analyze the case when $X$ is sparse CSparse . Further analysis of deterministic CholeskyQR-type algorithms can be found in Prob ; 2016 ; Mixed .
Algorithm 2 $[Q_{1},R_{2}]=\mbox{CholeskyQR2}(X)$
1:  $[Q,R]=\mbox{CholeskyQR}(X),$
2:  $[Q_{1},R_{1}]=\mbox{CholeskyQR}(Q),$
3:  $R_{2}=R_{1}R.$
Algorithm 3 $[Q,R]=\mbox{SCholeskyQR}(X)$
1:  $B=X^{\top}X,$
2:  choose $s>0$,
3:  $R=\mbox{Cholesky}(B+sI),$
4:  $Q=XR^{-1}.$
Algorithm 4 $[Q_{2},R_{4}]=\mbox{SCholeskyQR3}(X)$
1:  $[Q,R]=\mbox{SCholeskyQR}(X),$
2:  $[Q_{1},R_{1}]=\mbox{CholeskyQR}(Q),$
3:  $R_{2}=R_{1}R,$
4:  $[Q_{2},R_{3}]=\mbox{CholeskyQR}(Q_{1}),$
5:  $R_{4}=R_{3}R_{2}.$
1.0.2 LU-CholeskyQR2 and its bottleneck
Among all the deterministic CholeskyQR-type algorithms, LU-CholeskyQR LUChol is particularly noteworthy. It has the unique advantage of not imposing a restriction on $\kappa_{2}(X)$, which is crucial for real-world applications in industry. LU-CholeskyQR combines LU factorization with CholeskyQR, as shown in Algorithm 5. Here, $L\in\mathbb{R}^{m\times n}$ is a unit tall-skinny lower triangular matrix, and $U\in\mathbb{R}^{n\times n}$ is an upper triangular matrix when $X\in\mathbb{R}^{m\times n}$ is tall-skinny. A CholeskyQR step is performed afterwards to form LU-CholeskyQR2, ensuring orthogonality, as detailed in Algorithm 6.
Algorithm 5 $[Q,R]=\mbox{LU-CholeskyQR}(X)$
1:  $PA=LU,$
2:  $B=L^{\top}L,$
3:  $S=\mbox{Cholesky}(B),$
4:  $R=SU,$
5:  $Q=XR^{-1}.$
Algorithm 6 $[Q_{1},R_{2}]=\mbox{LU-CholeskyQR2}(X)$
1:  $[Q,R]=\mbox{LU-CholeskyQR}(X),$
2:  $[Q_{1},R_{1}]=\mbox{CholeskyQR}(Q),$
3:  $R_{2}=R_{1}R.$
In practice, the $L$-factor after LU factorization is often well-conditioned, even when $X$ is ill-conditioned. However, there are many real-world scenarios that can lead to an ill-conditioned $L$, such as certain lower-triangular matrices and diagonally dominant matrices with many negative elements. These situations frequently arise in linear equations, information theory, and systems theory. The primary bottleneck of LU-CholeskyQR2 is the potential for numerical breakdown in these cases, which is a significant problem that remains to be addressed or improved.
1.0.3 Randomized algorithms and our contributions
In recent years, randomized numerical linear algebra halko2011 ; 2020 has become a hot topic worldwide. A new randomized technique called sketching Ach has been widely applied to various problems. The sketching technique primarily aims to reduce computational costs by replacing the original large matrices with alternative matrices of smaller sizes after some preconditioning steps. The sketching matrix is typically chosen to be an $\epsilon$-subspace embedding or a linear map to a lower-dimensional space, preserving the inner products and norms of all vectors within the subspace up to a factor of $\sqrt{1+\epsilon}$, where $10\leq\epsilon<$ 1. Several existing methods for sketching matrices include Gaussian Sketch, SRHT, and Count Sketch rgs ; pmlr ; Fast ; Estimating ; pylspack . Research on randomized CholeskyQR has also emerged. Y. Fan and his collaborators Novel proposed the initial version of randomized CholeskyQR, while O. Balabanov Randomized provided a detailed analysis of CholeskyQR-type algorithms with various randomized strategies, including sketching. A recent work by A. J. Higgins and his collaborators Householder introduced a novel method called multi-sketching, which employs two different sketching steps consecutively. This work developed a new algorithm called randomized Householder QR, utilizing HouseholderQR to replace the original generation step of the $R$-factor in CholeskyQR to avoid numerical breakdown. The multi-sketching technique is designed to accelerate the entire algorithm. The corresponding algorithm is outlined in Algorithm 7. Here, $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ and $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ are sketching matrices, with $n\leq s_{2}\leq s_{1}\leq 1$. To ensure good orthogonality, a CholeskyQR step is added afterwards, resulting in the Rand.Householder-Cholesky algorithm, as shown in Algorithm 8. They provide a clear rounding error analysis of the algorithms with two $\epsilon$-subspace embeddings, and numerical experiments demonstrate that their randomized Householder-CholeskyQR is more efficient and numerically stable than the algorithms in Randomized .
Algorithm 7 $[Q,R]=\mbox{Randomized Householder QR}(X)$
1:  $X_{ss}=\Omega_{2}\Omega_{1}X,$
2:  $[Q_{ss},R]=\mbox{HouseholderQR}(X_{ss}),$
3:  $Q=XR^{-1}.$
Algorithm 8 $[Q_{1},R_{2}]=\mbox{Rand.Householder-Cholesky}(X)$
1:  $[Q,R]=\mbox{Randomized Householder QR}(X)$
2:  $[Q_{1},R_{1}]=\mbox{CholeskyQR}(Q),$
3:  $R_{2}=R_{1}R.$
In this work, we combine randomized Householder-CholeskyQR with LU-CholeskyQR2 to enhance the applicability of LU-CholeskyQR2. We construct a new deterministic method called LU-Householder CholeskyQR2 (LHC2) by partially replacing the generation of the $R$-factor with HouseholderQR. LHC2 can handle many cases that may lead to numerical breakdown in LU-CholeskyQR2, demonstrating better applicability compared to the original LU-CholeskyQR2. To accelerate LHC2, we employ both single-sketching and multi-sketching techniques, resulting in two new randomized algorithms: SLHC2 and SSLHC3. We perform rounding error analysis for all three new algorithms. Numerical examples show that both SLHC2 and SSLHC3 are more efficient than the deterministic LHC2 for large matrices, highlighting the advantages of randomized algorithms over deterministic ones. Compared to SLHC2, our SSLHC3, which utilizes multi-sketching, is more numerically stable and efficient. All algorithms exhibit good robustness, which is crucial for real-world applications. To the best of our knowledge, our work is the first to provide a detailed analysis of the improvements to LU-CholeskyQR2. Notably, all of our new algorithms do not require the sufficient condition of $\kappa_{2}(X)$, which is their most significant advantage compared to other CholeskyQR-type algorithms.
1.0.4 Outline of this work
Our work is organized as follows: We present existing results from previous studies in Section 2. In Sections 3 and 4, we introduce our improved deterministic and randomized algorithms, along with their corresponding rounding error analyses regarding orthogonality and residuals. The primary numerical experiments and comparisons are detailed in Section 5. Finally, we conclude with a summary of the results in Section 6.
2 Literature review
In this section, we present existing results from previous work related to rounding error analysis and the sketching technique.
2.1 Rounding error analysis
Here, we present some key theoretical results of rounding error analysis that are crucial to this work.
Lemma 1 (Weyl’s Theorem MatrixC )
For $A,B,C\in\mathbb{R}^{m\times n}$, if we have $A+B=C$, then
$|\sigma_{i}(A)-\sigma_{i}(B)|\leq\sigma_{1}(C)=\|C\|_{2}.$
where $\sigma_{i}(X)$ is the $i$-th greatest singular value of $X$, with $i=1,2,\cdots\min(m,n)$.
Lemma 2 (Rounding error in matrix multiplication Higham )
For $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times p}$, the error in computing the matrix product $C=AB$ in floating-point arithmetic is bounded by
$|AB-fl(AB)|\leq\gamma_{n}|A||B|.$
Here, $|A|$ is the matrix whose $(i,j)$ element is $|a_{ij}|$ and
$\gamma_{n}:=\frac{n{{\bf u}}}{1-n{{\bf u}}}\leq 1.02n{{\bf u}}.$
Lemma 3 (Rounding error in HouseholderQRHigham )
If an orthogonal factor $Q\in\mathbb{R}^{m\times n}$ and an upper-triangular factor $R\in\mathbb{R}^{n\times n}$ are received after doing a HouseholderQR of the input matrix $X\in\mathbb{R}^{m\times n}$,
$X+\Delta X=QR,\|\Delta X_{j}\|_{2}\leq\gamma_{mn}\|X_{j}\|_{2}.$
Here, $X_{j}$ denotes the $j$-th column of $X$, $j=1,2,\cdots,n$.
To learn more about the rounding error analysis of matrix perturbation, readers can refer to Improved ; Perturbation ; Backward for more details.
2.2 Matrix sketching
Matrix sketching is a widely used technique in various aspects of scientific computing. The general idea is to use random matrices to compress the dimensions of the original matrix while bounding the norm of the matrix and its columns. The purpose of matrix sketching is to represent the original matrix in a smaller and more economical form. In this section, we present some existing results related to matrix sketching.
2.2.1 Single-sketching
In this section, we introduce the concept of sketching with the definition of an $\epsilon$-subspace embedding rgs ; Fast ; 4031351 .
Definition 1 ($\epsilon$-subspace embedding)
When there is an $\epsilon$ which satisfies $10\leq\epsilon<$ 1, the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-subspace embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, if for any $x,y\in\mathcal{K}$,
$|\langle x,y\rangle-\langle\Omega x,\Omega y\rangle|\leq\epsilon\|x\|_{2}\|y\|_{2}.$ (1)
[…]
In the following, we list some properties based on Definition 1.
Lemma 4 (Single-sketching of the vector)
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, then for any $x\in\mathcal{K}$, we can have
$\sqrt{1-\epsilon}\cdot\|x\|_{2}\leq\|\Omega x\|_{2}\leq\sqrt{1+\epsilon}\cdot\|x\|_{2}.$ (2)
Lemma 4 focuses on the sketching of vectors. Below, we list the properties of the matrix after sketching.
Lemma 5 (Single-sketching of the matrix)
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, if $T$ is a matrix whose columns form a basis of $\mathcal{K}$, we can have
$\sqrt{1-\epsilon}\cdot\|T\|_{2}\leq\|\Omega T\|_{2}\leq\sqrt{1+\epsilon}\cdot\|\Omega T\|_{2},$ (3)
$\sqrt{1-\epsilon}\cdot\|T\|_{F}\leq\|\Omega T\|_{F}\leq\sqrt{1+\epsilon}\cdot\|\Omega T\|_{F},$ (4)
$\sqrt{1-\epsilon}\cdot\sigma_{min}(T)\leq\sigma_{min}(\Omega T)\leq\|\Omega T\|_{2}\leq\sqrt{1+\epsilon}\cdot\|T\|_{2}.$ (5)
Here, $\|T\|_{2}=\sigma_{1}(T)$ denotes the biggest singular value of $T$.
In Lemma 5, equations (3) and (5) play a crucial role in our analysis. Equation (4) can be easily derived from (3). Below, we list some existing results related to condition numbers.
Lemma 6 (Condition number of the single-sketching)
When the sketch matrix $\Omega\in\mathcal{R}^{s\times n}$ is an $\epsilon$-subspace embedding for the subspace $\mathcal{K}\subset\mathbb{R}^{n}$, if $T$ is a matrix whose columns form a basis of $\mathcal{K}$, we can have
$\frac{\sigma_{min}(\Omega T)}{\sqrt{1+\epsilon}}\leq\sigma_{min}(T)\leq\|T\|_{2}\leq\frac{\|\Omega T\|_{2}}{\sqrt{1-\epsilon}}.$ (6)
That is,
$\kappa_{2}(T)\leq\sqrt{\frac{1-\epsilon}{1+\epsilon}}\cdot\kappa_{2}(\Omega T).$ (8)
Here, $\kappa_{2}(T)=\frac{\|T\|_{2}}{\sigma_{min}(T)}$ denotes the condition number of $T$.
The results mentioned above pertain to $\epsilon$-subspace embeddings, which require knowledge of the subspace $\mathcal{K}$. In contrast, a concept known as the $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding rgs does not require this knowledge.
Definition 2 ($(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding)
When $\Omega\in\mathbb{R}^{s\times m}$ is an $\epsilon$-subspace embedding for all the fixed $n$-dimensional subspace $\mathcal{K}\subset\mathbb{R}^{m}$ with probability at least $1-p$, it is an $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding.
Typical $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embeddings include both dense and sparse cases, which correspond to different sketching methods. A sketching matrix is defined as $\Omega=\frac{1}{s}G$, where $G\in\mathbb{R}^{s\times m}$ is a Gaussian matrix. The sketching size $s$ 4031351 satisfies
$s=\eta\cdot\frac{logn\cdot log{\frac{1}{p}}}{\epsilon^{2}}.$ (9)
Here, $\eta$ is a positive constant. In practical applications, we can set $s=kn$, where $k$ is a small positive constant.
Regarding $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embeddings, researchers often use CountSketch Low . If $\Omega\in\mathbb{R}^{s\times m}$ is a sketching matrix for CountSketch, then, according to frequent ; tool , $s$ needs to satisfy
$s\geq\frac{n^{2}+n}{\epsilon^{2}p}.$ (10)
2.2.2 Multi-sketching
In Householder , a technique called multi-sketching is introduced, which involves applying two steps of sketching consecutively. Below, we list some properties of the multi-sketching technique.
Lemma 7 (Multi-sketching of the vector)
If $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ is a $(\epsilon_{2},p_{2},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$ and they are generated independently, then for any $n$-dimensional subspace $\mathcal{K}\subset\mathbb{R}^{m}$ and any $x\in\mathcal{K}$, we can have
$\sqrt{1-\epsilon_{s}}\cdot\|x\|_{2}\leq\|\Omega_{2}\Omega_{1}x\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|x\|_{2}$ (11)
with probability at least $1-p_{f}$. Here, $p_{f}=p_{1}+p_{2}-p_{1}p_{2}$, $\epsilon_{s}=\epsilon_{1}+\epsilon_{2}-\epsilon_{1}\epsilon_{2}$ and $\epsilon_{b}=\epsilon_{1}+\epsilon_{2}+\epsilon_{1}\epsilon_{2}$.
Moreover, similar to Lemma 5 and Lemma 6, we can derive the following results.
Lemma 8 (Multi-sketching of the matrix)
With similar settings in Lemma 7, for $T\in\mathbb{R}^{m\times n}$, we can have
$\sqrt{1-\epsilon_{s}}\cdot\|T\|_{2}\leq\|\Omega_{2}\Omega_{1}T\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{2},$ (12)
$\sqrt{1-\epsilon_{s}}\cdot\|T\|_{F}\leq\|\Omega_{2}\Omega_{1}T\|_{F}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{F},$ (13)
$\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(T)\leq\sigma_{min}(\Omega_{2}\Omega_{1}T)\leq\|\Omega_{2}\Omega_{1}T\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{2},$ (14)
$\frac{\sigma_{min}(\Omega T)}{\sqrt{1+\epsilon_{b}}}\leq\sigma_{min}(T)\leq\|T\|_{2}\leq\frac{\|\Omega T\|_{2}}{\sqrt{1-\epsilon_{s}}}$ (15)
with probability at least $1-p_{f}$. With (15), we can have
$\kappa_{2}(T)\leq\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}\cdot\kappa_{2}(\Omega T)$ (16)
with probability at least $1-p_{f}$.
Among all sketching techniques, CountSketch offers low computational and storage costs. However, as indicated by (10), it requires a relatively large sketching size $s$ to maintain the $\epsilon$-embedding properties, which can result in significantly higher FLOPs after HouseholderQR. In contrast, the Gaussian sketch requires a much smaller sketching size than CountSketch, as shown in (9). However, it occupies more space and takes longer to sketch the dense matrix $X$. By applying a Gaussian sketch after a CountSketch, we can achieve low computational and storage costs while leaving a smaller matrix for HouseholderQR, thereby enhancing the algorithm’s efficiency. This is the primary advantage of multi-sketching over traditional single sketching. In this work, we focus on multi-sketching with a Gaussian sketch following a CountSketch. The theoretical analysis steps for other types of sketches are similar and can be explored by the readers.
3 Deterministic LU-Householder CholeskyQR
In this section, we introduce a preconditioning step known as LU-Householder CholeskyQR (LHC), which differs from LU-CholeskyQR. LHC is formed by combining LU-CholeskyQR and HouseholderQR. The original steps for calculating the Gram matrix and performing Cholesky factorization are replaced with HouseholderQR to avoid numerical breakdown when the input $X$ has an ill-conditioned $L$ after LU factorization. This process is detailed in Algorithm 9. To ensure numerical stability with respect to both orthogonality and residuals, an additional step of CholeskyQR is applied after LHC, resulting in LHC2, as presented in Algorithm 10.
Algorithm 9 $[Q,R]=\mbox{LHC}(X)$
1:  $PX=LU,$
2:  $[W,G]=\mbox{HouseholderQR}(L),$
3:  $R=GU,$
4:  $Q=XR^{-1}.$
Algorithm 10 $[Q_{1},R_{2}]=\mbox{LHC2}(X)$
1:  $[Q,R]=\mbox{LHC}(X),$
2:  $[Q_{1},R_{1}]=\mbox{CholeskyQR}(Q),$
3:  $R_{2}=R_{1}R.$
3.1 Theoretical results of LHC2 and general settings
Before providing theoretical results of LHC2, we write LHC first with error matrices as below:
$PX$ $=$ $LU+E_{LU},$ (17)
$L$ $=$ $WG+E_{HQ},$ (18)
$R$ $=$ $GU-E_{R},$ (19)
$X$ $=$ $QR-E_{QR}.$ (20)
If $X\in\mathbb{R}^{m\times n}$, we make the following assumptions:
$mn\sqrt{n}{\bf u}$$\leq$ $\frac{1}{64},$ (21)
$n^{2}{\bf u}$$\leq$ $\frac{1}{64},$ (22)
$\kappa_{2}(L)\sqrt{mn\sqrt{n}{\bf u}}$$\leq$ $1,$ (23)
$64\kappa_{2}(L)\kappa_{2}(U)n^{2}{\bf u}$$\leq$ 1.1 1.1 . (24)
Here, ${\bf u}$ is the machine precision and ${\bf u}=2^{-53}$.
In the following, we provide the theoretical results of rounding error analysis of LHC2.
Theorem 3.1 (Rounding error analysis of LHC2)
If $[Q_{1},R_{2}]=LHC2(X)$, $X\in\mathbb{R}^{m\times n}$, based on our previous settings, we can have
$\|Q_{1}^{\top}Q_{1}-I\|_{F}$$\leq$ $34.85(mn{\bf u}+n(n+1){\bf u}),$ (25)
$\|Q_{1}R_{2}-X\|_{F}$$\leq$ $7.94n^{2}{\bf u}\|X\|_{2}.$ (26)
3.2 Preparations for proving Theorem 3.1
Before proving this theorem, we do some preparations and analysis first. We focus on LHC first. With (20), we can have
$Q^{\top}Q$ $=$ $R^{-\top}(X+E_{QR})^{\top}(X+E_{QR})R^{-1}$ (27)
$=$ $R^{-\top}(X^{\top}X+X^{\top}E_{QR}+E_{QR}^{\top}X+E_{QR}^{\top}E_{QR})R^{-1}$
$=$ $R^{-\top}X^{\top}XR^{-1}+R^{-\top}(X^{\top}E_{QR}+E_{QR}^{\top}X+E_{QR}^{\top}E_{QR})R^{-1}.$
With (27), we can further decompose $R^{-\top}X^{\top}XR^{-1}$. Based on (17), we can have
$R^{-\top}X^{\top}XR^{-1}$ $=$ $R^{-\top}(LU+E_{LU})^{\top}(LU+E_{LU})R^{-1}$ (28)
$=$ $R^{-\top}U^{\top}L^{\top}LUR^{-1}+R^{-\top}(E_{LU}^{\top}LU+U^{\top}L^{\top}E_{LU}+E_{LU}^{\top}E_{LU})R^{-1}.$
We try to transform $R^{-\top}U^{\top}L^{\top}LUR^{-1}$ in (28). With (18) and (19), when $W$ is an orthogonal factor, we can get
$R^{-\top}U^{\top}L^{\top}LUR^{-1}$ $=$ $R^{-\top}U^{\top}(WG+E_{HQ})^{\top}(WG+E_{HQ})UR^{-1}$ (29)
$=$ $R^{-\top}U^{\top}G^{\top}W^{\top}WGUR^{-1}$
$+$ $R^{-\top}U^{\top}(E_{HQ}^{\top}E_{HQ}+G^{\top}W^{\top}E_{HQ}+E_{HQ}^{\top}WG)UR^{-1}$
$=$ $R^{-\top}(R+E_{R})^{\top}(R+E_{R})R^{-1}$
$+$ $R^{-\top}U^{\top}(E_{HQ}^{\top}E_{HQ}+G^{\top}W^{\top}E_{HQ}+E_{HQ}^{\top}WG)UR^{-1}$
$=$ $R^{-\top}(R^{\top}R+E_{R}^{\top}R+R^{\top}E_{R}+E_{R}^{\top}E_{R})R^{-1}$
$+$ $R^{-\top}U^{\top}(E_{HQ}^{\top}E_{HQ}+G^{\top}W^{\top}E_{HQ}+E_{HQ}^{\top}WG)UR^{-1}$
$=$ $I+(R^{-\top}E_{R}^{\top}+E_{R}R^{-1}+R^{-\top}E_{R}^{\top}E_{R}R^{-1})$
$+$ $R^{-\top}U^{\top}(E_{HQ}^{\top}E_{HQ}+G^{\top}W^{\top}E_{HQ}+E_{HQ}^{\top}WG)UR^{-1}.$
With (27)-(29), we can set
$\|Q^{\top}Q-I\|_{2}\leq\|E_{1}\|_{2}+\|E_{2}\|_{2}+\|E_{3}\|_{2}$ (30)
where
$E_{1}$ $=$ $R^{-\top}(X^{\top}E_{QR}+E_{QR}^{\top}X+E_{QR}^{\top}E_{QR})R^{-1},$ (31)
$E_{2}$ $=$ $R^{-\top}U^{\top}(E_{HQ}^{\top}E_{HQ}+G^{\top}W^{\top}E_{HQ}+E_{HQ}^{\top}WG)UR^{-1}$ (32)
$+$ $R^{-\top}(E_{LU}^{\top}LU+U^{\top}L^{\top}E_{LU}+E_{LU}^{\top}E_{LU})R^{-1},$
$E_{3}$ $=$ $(R^{-\top}E_{R}^{\top}+E_{R}R^{-1}+R^{-\top}E_{R}^{\top}E_{R}R^{-1}).$ (33)
With (31)-(33), we can have
$\|E_{1}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{QR}\|_{2}\|XR^{-1}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{QR}\|_{2}^{2},$ (34)
$\|E_{2}\|_{2}$$\leq$ $\|UR^{-1}\|_{2}^{2}\|E_{HQ}\|_{2}^{2}+2\|UR^{-1}\|_{2}\|GUR^{-1}\|_{2}\|E_{HQ}\|_{2}$ (35)
$+$ $2\|E_{LU}\|_{2}\|R^{-1}\|_{2}\|LUR^{-1}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{LU}\|_{2}^{2},$
$\|E_{3}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{R}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{R}\|_{2}^{2}.$ (36)
Therefore, with (34)-(36), if we want to estimate (30), we need to estimate $\|R^{-1}\|_{2}$, $\|E_{QR}\|_{2}$, $\|XR^{-1}\|_{2}$, $\|UR^{-1}\|_{2}$, $\|E_{HQ}\|_{2}$, $\|GUR^{-1}\|_{2}$, $\|LUR^{-1}\|_{2}$ and $\|E_{R}\|_{2}$. We do the corresponding estimations below.
Lemma 9 (Estimation of $\|E_{HQ}\|_{2}$)
In (18), we can have
$\|E_{HQ}\|_{2}\leq 1.1mn\sqrt{n}{\bf u}\|L\|_{2}.$ (37)
Proof
According to Lemma 3 and (18), when $W\in\mathbb{R}^{m\times n}$ is orthogonal and $G\in\mathbb{R}^{n\times n}$, we can have
$\|E_{HQj}\|_{2}\leq\gamma_{mn}\|L_{j}\|_{2}.$ (38)
Here, $E_{HQj}$ and $L_{j}$ denote the $j$-th column of $E_{HQ}$ and $L$, $j=1,2,\cdots,n$. We sum all the $n$ inequalities together and we can get
$\|E_{HQ}\|_{2}$$\leq$ $\|E_{HQ}\|_{F}$
$\leq$ $\gamma_{mn}\|L\|_{F}$
$\leq$ $\gamma_{mn}\cdot\sqrt{n}\|L\|_{2}$
$\leq$ $1.1mn\sqrt{n}{\bf u}\|L\|_{2}.$
(37) is proved.
Lemma 10 (Estimation of $\|E_{R}\|_{2}$)
In (19), we can have
$\|E_{R}\|_{2}\leq 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}.$ (39)
Proof
According to Lemma 2 and (19), we can have
$\|E_{R}\|_{2}\leq 1.1n^{2}{\bf u}\|G\|_{2}\|U\|_{2}.$ (40)
For $\|G\|_{2}$, with (18) and (21), when $W$ is orthogonal, we can have
$\|G\|_{2}$ $=$ $\|WG\|_{2}$ (41)
$\leq$ $\|L\|_{2}+\|E_{HQ}\|_{2}$
$\leq$ $(1+1.1mn\sqrt{n}{\bf u})\|L\|_{2}$
$\leq$ $1.02\|L\|_{2}.$
We put (41) into (40) and we can get
$\|E_{R}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|U\|_{2}\cdot 1.02\|L\|_{2}$
$\leq$ $1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}.$
(39) is proved.
Lemma 11 (Estimation of $\|R^{-1}\|_{2}$)
For $\|R^{-1}\|_{2}$, we can have
$\|R^{-1}\|_{2}\leq\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}.$ (42)
Proof
Based on Lemma 1 and (19), we can have
$\sigma_{min}(R)$ $=$ $\sigma_{min}(GU-E_{R})$ (43)
$\geq$ $\sigma_{min}(G)\sigma_{min}(U)-\|E_{R}\|_{2}.$
For $\sigma_{min}(G)$, with Lemma 1, (18), (21), (23) and (37), we can have
$\sigma_{min}(G)$ $=$ $\sigma_{min}(WG)$ (44)
$\geq$ $\sigma_{min}(L)-\|E_{HQ}\|_{2}$
$\geq$ $\sigma_{min}(L)-1.1mn\sqrt{n}{\bf u}\|L\|_{2}$
$\geq$ $0.8\sigma_{min}(L).$
We put (44) into (43) and combine them with (21)-(24). Therefore, we can get
$\sigma_{min}(R)$ $=$ $0.8\sigma_{min}(L)\sigma_{min}(U)-1,13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}$ (45)
$\geq$ $0.8\sigma_{min}(L)\sigma_{min}(U)-\frac{1.13}{64}\cdot\sigma_{min}(L)\sigma_{min}(U)$
$\geq$ $0.782\sigma_{min}(L)\sigma_{min}(U).$
With (45), we can easily receive (42). (42) is proved.
Lemma 12 (Estimation of $\|UR^{-1}\|_{2}$)
For $\|UR^{-1}\|_{2}$, we can have
$\|UR^{-1}\|_{2}\leq\frac{1.29}{\sigma_{min}(L)}.$ (46)
Proof
We multiply both sides of (19) with $R^{-1}$ and we can get
$I+E_{R}R^{-1}=GUR^{-1}.$ (47)
Here, $I\in\mathbb{R}^{n\times n}$ is the identity matrix. We multiply both sides of (47) with $G^{-1}$ and take norms. Thus, we can receive
$\|UR^{-1}\|_{2}=\|G^{-1}\|_{2}(1+\|E_{R}\|_{2}\|R^{-1}\|_{2}).$ (48)
With (44), we can get
$\|G^{-1}\|_{2}\leq\frac{1.25}{\sigma_{min}(L)}.$ (49)
We put (39), (42) and (49) into (48) and with (24), we can get
$\|UR^{-1}\|_{2}$$\leq$ $\|G^{-1}\|_{2}(1+\|E_{R}\|_{2}\|R^{-1}\|_{2})$
$\leq$ $\frac{1.25}{\sigma_{min}(L)}\cdot(1+1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})$
$\leq$ $\frac{1.25}{\sigma_{min}(L)}\cdot(1+1.45n^{2}{\bf u}\kappa_{2}(L)\kappa_{2}(U))$
$\leq$ $\frac{1.29}{\sigma_{min}(L)}.$
(46) is proved.
Lemma 13 (Estimation of $\|LUR^{-1}\|_{2}$)
For $\|LUR^{-1}\|_{2}$, we can have
$\|LUR^{-1}\|_{2}\leq 2.$ (50)
Proof
With (18) and (19), when $W$ is orthogonal, we can get
$\|LUR^{-1}\|_{2}^{2}$ $=$ $\|R^{-\top}U^{\top}L^{\top}LUR^{-1}\|_{2}$ (51)
$=$ $\|R^{-\top}U^{\top}(WG+E_{HQ})^{\top}(WG+E_{HQ})UR^{-1}\|_{2}$
$\leq$ $\|R^{-\top}U^{\top}G^{\top}GUR^{-1}\|_{2}+\|UR^{-1}\|_{2}^{2}(\|E_{HQ}\|_{2}^{2}+2\|G\|_{2}\|E_{HQ}\|_{2})$
$=$ $\|R^{-1}(R+E_{R})\|_{2}^{2}+\|UR^{-1}\|_{2}^{2}(\|E_{HQ}\|_{2}^{2}+2\|G\|_{2}\|E_{HQ}\|_{2})$
$\leq$ $(1+\|E_{R}\|_{2}\|R^{-1}\|_{2})^{2}+\|UR^{-1}\|_{2}^{2}(\|E_{HQ}\|_{2}^{2}+2\|G\|_{2}\|E_{HQ}\|_{2}).$
Here, we put (21), (23), (24), (37), (39), (41), (42) and (46) into (51) and we can get
$\|LUR^{-1}\|_{2}^{2}$$\leq$ $(1+\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}$ (52)
$+$ $(\frac{1.29}{\sigma_{min}(L)})^{2}\cdot(1.21m^{2}n^{3}{\bf u}^{2}\|L\|_{2}^{2}+2\cdot 1.02\|L\|_{2}\cdot 1.1mn\sqrt{n}{\bf u}\|L\|_{2})$
$\leq$ $(1+0.03)^{2}+(\frac{1.29}{\sigma_{min}(L)})^{2}\cdot(1.21\cdot\frac{1}{64}\cdot\sigma_{min}(L)^{2}+2\cdot 1.02\cdot 1.1\sigma_{min}(L)^{2})$
$\leq$ $1.07+2.90$
$\leq$ $4$
With (52), (50) is proved.
Lemma 14 (Estimation of $\|GUR^{-1}\|_{2}$)
For $\|GUR^{-1}\|_{2}$, we can have
$\|GUR^{-1}\|_{2}\leq 1.03.$ (53)
Proof
Based on (47), with (24), (39) and (42), we can get
$\|GUR^{-1}\|_{2}$$\leq$ $\|I\|_{2}+\|E_{R}\|_{2}\|R^{-1}\|_{2}$
$\leq$ $1+1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}$
$\leq$ $1.03$
(53) is proved.
Lemma 15 (Estimation of $\|XR^{-1}\|_{2}$)
For $\|XR^{-1}\|_{2}$, we can have
$\|XR^{-1}\|_{2}\leq 2.02.$ (54)
Proof
With (17), we can have
$\|XR^{-1}\|_{2}^{2}$ $=$ $\|R^{-\top}X^{\top}XR^{-1}\|_{2}$ (55)
$=$ $\|R^{-\top}U^{\top}L^{\top}LUR^{-1}+R^{-\top}(E_{LU}^{\top}LU+U^{\top}L^{\top}E_{LU}+E_{LU}^{\top}E_{LU})R^{-1}\|_{2}$
$\leq$ $\|LUR^{-1}\|_{2}^{2}+2\|E_{LU}\|_{2}\|R^{-1}\|_{2}\|LUR^{-1}\|_{2}+\|E_{LU}\|_{2}^{2}\|R^{-1}\|_{2}^{2}.$
To estimate $\|E_{LU}\|_{2}$, with Lemma 2 and (17), we can have
$\|E_{LU}\|_{2}\leq 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}.$ (56)
We put (42), (50) and (56) into (55) and with (24), we can have
$\|XR^{-1}\|_{2}^{2}$$\leq$ $\|LUR^{-1}\|_{2}^{2}+2\|E_{LU}\|_{2}\|R^{-1}\|_{2}\|LUR^{-1}\|_{2}+\|E_{LU}\|_{2}^{2}\|R^{-1}\|_{2}^{2}$ (57)
$\leq$ $3.97+2\cdot 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2$
$+$ $(1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}\cdot(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})^{2}$
$\leq$ $3.97+2\cdot\frac{1.1\cdot 1.28\cdot 2}{64}+(\frac{1.1\cdot 1.28}{64})^{2}$
$\leq$ $4.06$
With (57), we can have (54). (54) is proved.
Lemma 16 (Estimation of $\|E_{QR}\|_{2}$)
For $\|E_{QR}\|_{2}$, we can have
$\|E_{QR}\|_{2}\leq 2.33n^{2}{\bf u}\|R\|_{2}.$ (58)
Proof
With Lemma2, (20), we can have
$\|E_{QR}\|_{2}\leq 1.1n^{2}\|Q\|_{2}\|R\|_{2}.$ (59)
For $\|Q\|_{2}$, based on Lemma 2 and (20), we can have
$\|Q\|_{2}$$\leq$ $\|XR^{-1}\|_{2}+\|E_{QR}R^{-1}\|_{2}$ (60)
$\leq$ $\|XR^{-1}\|_{2}+1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}\cdot\|R^{-1}\|_{2}$
$\leq$ $\|XR^{-1}\|_{2}+1.1n^{2}{\bf u}\kappa_{2}(R)\|Q\|_{2}.$
For $\kappa_{2}(R)$, we need to estimate $\|R\|_{2}$ first. With (19), (22), (39) and (41), we can get
$\|R\|_{2}$$\leq$ $\|G\|_{2}\|U\|_{2}+\|E_{R}\|_{2}$ (61)
$\leq$ $1.02\|L\|_{2}\|U\|_{2}+1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}$
$\leq$ $1.04\|L\|_{2}\|U\|_{2}.$
We combine (42) with (61) and we can estimate $\kappa_{2}(R)$ as
$\kappa_{2}(R)$ $=$ $\|R\|_{2}\|R^{-1}\|_{2}$ (62)
$\leq$ $1.04\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}$
$\leq$ $1.34\kappa_{2}(L)\kappa_{2}(U).$
We put (54) and (62) into (63) and with (24), we can have
$\|Q\|_{2}$$\leq$ $\frac{\|XR^{-1}\|_{2}}{1-1.1n^{2}{\bf u}\kappa_{2}(R)}$ (63)
$\leq$ $\frac{2.02}{1-1.1\cdot 1.34\cdot\frac{1}{64}}$
$\leq$ $2.07$
We put (62) and (63) into (59) and we can have (58). (58) is proved.
3.3 Proof of Theorem 3.1
With all the lemmas above, we start to prove Theorem 3.1. Some steps and ideas are similar to that of LUChol but there are also many differences.
Proof
Our proof of Theorem 3.1 has two parts, orthogonality and residual.
3.3.1 Orthogonality
We estimate the orthogonality of LHC2 first. According to (30), we estimate $\|E_{1}\|_{2}$, $\|E_{2}\|_{2}$ and $\|E_{3}\|_{2}$ seperately.
For $\|E_{1}\|_{2}$, we put (42), (54) and (58) into (34). With (24) and (61), we can get
$\|E_{1}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{QR}\|_{2}\|XR^{-1}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{QR}\|_{2}^{2}$ (64)
$\leq$ $2\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2.33n^{2}{\bf u}\|R\|_{2}\cdot 2.02$
$+$ $(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})^{2}\cdot(2.33n^{2}{\bf u}\|R\|_{2})^{2}$
$\leq$ $2\cdot 1.28\cdot 1.04\cdot 2.33n^{2}{\bf u}\kappa_{2}(L)\kappa_{2}(U)\cdot 2.02$
$+$ $(1.28\cdot 2.33\cdot 1.04)^{2}\cdot(n^{2}{\bf u}\kappa_{2}(L)\kappa_{2}(U))^{2}$
$\leq$ $2\cdot 1.28\cdot 1.04\cdot 2.03\cdot 2.02\cdot\frac{1}{64}+(1.28\cdot 2.33\cdot 1.04)^{2}\cdot(\frac{1}{64})^{2}$
$\leq$ $0.2$
For $\|E_{2}\|_{2}$, we put (37), (42), (46), (50), (53) and (56) into (35). With (21), (23) and (24), we can have
$\|E_{2}\|_{2}$$\leq$ $\|UR^{-1}\|_{2}^{2}\|E_{HQ}\|_{2}^{2}+2\|UR^{-1}\|_{2}\|GUR^{-1}\|_{2}\|E_{HQ}\|_{2}$ (65)
$+$ $2\|E_{LU}\|_{2}\|R^{-1}\|_{2}\|LUR^{-1}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{LU}\|_{2}^{2}$
$\leq$ $(\frac{1.29}{\sigma_{min}(L)}\cdot 1.1mn\sqrt{n}{\bf u}\|L\|_{2})^{2}+2\cdot\frac{1.29}{\sigma_{min}(L)}\cdot 1.03\cdot 1.1mn\sqrt{n}{\bf u}\|L\|_{2}$
$+$ $2\cdot 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2+(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}$
$\leq$ $(1.29\cdot 1.1\cdot\frac{1}{8})^{2}+2\cdot 1.29\cdot 1.03\cdot 1.1\cdot\frac{1}{8}+2\cdot 2\cdot 1.28\cdot 1.1\cdot\frac{1}{64}+(1.28\cdot 1.1\cdot\frac{1}{64})^{2}$
$\leq$ $0.49$
For $\|E_{3}\|_{2}$, we put (39) and (42) into (36). With (24), we can get
$\|E_{3}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{R}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{R}\|_{2}^{2}$ (66)
$\leq$ $2\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}+(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}$
$\leq$ $2\cdot 1.28\cdot 1.13\cdot\frac{1}{64}+(1.28\cdot 1.13\cdot\frac{1}{64})^{2}$
$\leq$ $0.05$
Therefore, we put (64)-(66) into (30) and we can get
$\|Q^{\top}Q-I\|_{2}\leq 0.74.$ (67)
With (67), we can get
$\|Q\|_{2}$$\leq$ $\sqrt{1+0.74}$ (68)
$\leq$ $1.32,$
$\sigma_{min}(Q)$$\geq$ $\sqrt{1-0.74}$ (69)
$\geq$ $0.5$
Based on (68) and (69), we can get
$\kappa_{2}(Q)$$\leq$ $\frac{\|Q\|_{2}}{\sigma_{min}(Q)}$ (70)
$\leq$ $2.64$
According to the theoretical result in error , for LHC2 with a step of CholeskyQR after LHC, based on (70), we can get
$\|Q_{1}^{\top}Q_{1}-I\|_{F}$$\leq$ $\frac{5}{64}\cdot(8\kappa_{2}(Q)\sqrt{mn{\bf u}+n(n+1){\bf u}})^{2}$
$\leq$ $5\cdot(\kappa_{2}(Q))^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$
$\leq$ $34.85(mn{\bf u}+n(n+1){\bf u}).$
(25) is proved.
3.3.2 Residual
In the following, we consider the residual. We write CholeskyQR after LHC with error matrices below.
$B=Q^{\top}Q+E_{A},$ (71)
$R_{1}^{\top}R_{1}=B+E_{B},$ (72)
$Q_{1}R_{1}=Q+E_{C},$ (73)
$R_{2}=R_{1}R+E_{D}.$ (74)
Based on (73) and (74), we can have
$Q_{1}R_{2}$ $=$ $(Q+E_{C})R_{1}^{-1}(R_{1}R+E_{D})$ (75)
$=$ $(Q+E_{C})R+Q_{1}E_{D}$
$\leq$ $QR+E_{C}R+Q_{1}E_{D}.$
With (75), we can get
$\|Q_{1}R_{2}-X\|_{F}\leq\|QR-X\|_{F}+\|E_{C}\|_{F}\|R\|_{2}+\|Q_{1}\|_{2}\|E_{D}\|_{F}.$ (76)
With Lemma 2, we can have
$\|E_{C}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|Q_{1}\|_{2}\|R_{1}\|_{2},$ (77)
$\|E_{D}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|R_{1}\|_{2}\|R\|_{2},$ (78)
$\|QR-X\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}.$ (79)
We put (77)-(79) into (76) and we can get
$\|QR-X\|_{F}$$\leq$ $\|QR-X\|_{F}+\|E_{C}\|_{F}\|R\|_{2}+\|Q_{1}\|_{2}\|E_{D}\|_{F}$ (80)
$\leq$ $1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}+1.1n^{2}{\bf u}\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}+1.1n^{2}{\bf u}\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}$
$\leq$ $1.1n^{2}{\bf u}(\|Q\|_{2}\|R\|_{2}+2\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2}).$
With (21), (22), (25) and the theoretical result in error , we can get
$\|Q_{1}\|_{2}$$\leq$ $\sqrt{1+34.85\cdot(\frac{1}{64}+\frac{1}{64})}$ (81)
$\leq$ $1.45,$
$\|R_{1}\|_{2}$$\leq$ $1.1\|Q\|_{2}$ (82)
$\leq$ $1.1\cdot 1.32$
$\leq$ $1.46$
With Lemma 1, (20), (79) and properties of singular values, we can have
$\|X\|_{2}$ $=$ $\|QR-E_{QR}\|_{2}$ (83)
$\geq$ $\sigma_{min}(QR-E_{QR})$
$\geq$ $\sigma_{min}(Q)\|R\|_{2}-1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}.$
With (22), (68), (69) and (83), we can have
$\|R\|_{2}$$\leq$ $\frac{\|X\|_{2}}{\sigma_{min}(Q)-1.1n^{2}{\bf u}\|Q\|_{2}}$ (84)
$\leq$ $\frac{\|X\|_{2}}{0.5-1.1\cdot\frac{1}{64}\cdot 1.32}$
$\leq$ $2.1\|X\|_{2}.$
Therefore, we put (68), (81), (82) and (84) into (80) and we can have
$\|QR-X\|_{F}$$\leq$ $1.1n^{2}{\bf u}(\|Q\|_{2}\|R\|_{2}+2\|Q_{1}\|_{2}\|R_{1}\|_{2}\|R\|_{2})$
$\leq$ $1.1n^{2}{\bf u}\cdot(1.32\cdot 2.1\|X\|_{2}+1.45\cdot 1.46\cdot 2.1\|X\|_{2})$
$\leq$ $7.94n^{2}{\bf u}\|X\|_{2}.$
(26) is proved. In all, Theorem 3.1 is proved.
4 Randomized LU-Householder CholeskyQR
In this section, we introduce the randomized versions of LU-Householder CholeskyQR (LHC): SLHC, which utilizes single-sketching, and SSLHC, which employs multi-sketching. These methods are detailed in Algorithm 11 and Algorithm 12, respectively. The single-sketching and multi-sketching techniques introduced in Section 4 are used to accelerate HouseholderQR in LHC. Compared to LHC2, SLHC includes a single step of CholeskyQR, resulting in SLHC2, as shown in Algorithm 13. For SSLHC, to achieve better orthogonality, we add CholeskyQR2 afterward, forming SSLHC3, as detailed in Algorithm 14. We also provide a theoretical analysis of the two algorithms, SLHC2 and SSLHC3, in this section.
Algorithm 11 $[Q_{s},R_{s}]=\mbox{SLHC}(X)$
1:  $PX=LU,$
2:  $L_{s}=\Omega_{1}L,$
3:  $[Q_{ts},G_{s}]=\mbox{Householder}(L_{s}),$
4:  $R_{s}=G_{s}U,$
5:  $Q_{s}=XR_{s}^{-1}.$
Algorithm 12 $[Q_{ss},R_{ss}]=\mbox{SSLHC}(X)$
1:  $PX=LU,$
2:  $L_{ss}=\Omega_{2}\Omega_{1}L,$
3:  $[Q_{tss},G_{ss}]=\mbox{Householder}(L_{ss}),$
4:  $R_{ss}=G_{ss}U,$
5:  $Q_{ss}=XR_{ss}^{-1}.$
Algorithm 13 $[Q_{s1},R_{s2}]=\mbox{SLHC2}(X)$
1:  $[Q_{s},R_{s}]=\mbox{SLHC}(X),$
2:  $[Q_{s1},R_{s1}=\mbox{CholeskyQR}(X),$
3:  $R_{s2}=R_{s1}R_{s}.$
Algorithm 14 $[Q_{ss2},R_{ss4}]=\mbox{SSLHC3}(X)$
1:  $[Q_{ss},R_{ss}]=\mbox{SSLHC}(X),$
2:  $[Q_{ss1},R_{ss1}=\mbox{CholeskyQR}(Q_{ss}),$
3:  $R_{ss2}=R_{ss1}R_{ss},$
4:  $[Q_{ss2},R_{ss3}=\mbox{CholeskyQR}(Q_{ss1}),$
5:  $R_{ss4}=R_{ss3}R_{ss2}.$
4.1 Theoretical results of SSLHC3 and general settings
In this part, we analyze SSLHC3. We write SSLHC first with error matrices below:
$PX$ $=$ $LU+E_{LU},$ (85)
$L_{ss}$ $=$ $\Omega_{2}\Omega_{1}L+E_{Lss},$ (86)
$L_{ss}$ $=$ $W_{ss}G_{ss}+E_{HQss},$ (87)
$R_{ss}$ $=$ $G_{ss}U-E_{Rss},$ (88)
$X$ $=$ $Q_{ss}R_{ss}-E_{QRss},$ (89)
$B_{1ss}$ $=$ $Q_{ss}^{\top}Q_{ss}+E_{1ss},$ (90)
$R_{1ss}^{\top}R_{1ss}$ $=$ $B_{1ss}+E_{2ss},$ (91)
$Q_{1ss}R_{1ss}$ $=$ $Q_{ss}+E_{3ss},$ (92)
$R_{2ss}$ $=$ $R_{1ss}R_{ss}+E_{4ss},$ (93)
$B_{2ss}$ $=$ $Q_{1ss}^{\top}{Q_{2ss}}+E_{5ss},$ (94)
$R_{3ss}^{\top}R_{3ss}$ $=$ $B_{2ss}+E_{6ss},$ (95)
$Q_{2ss}R_{3ss}$ $=$ $Q_{1ss}+E_{7ss},$ (96)
$R_{4ss}$ $=$ $R_{3ss}R_{2ss}+E_{8ss}.$ (97)
Here, $X\in\mathbb{R}^{m\times n}$, $S_{1}\in\mathbb{R}^{s_{1}\times m}$, $S_{2}\in\mathbb{R}^{s_{2}\times S_{1}}$, $n\leq s_{2}\leq s_{1}\leq m$.
To analyze SSLHC with the multi-sketching, we focus on $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding. We provide the following assumption.
Assumption 1 (The assumption for the multi-sketching)
If $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, $\Omega_{2}\in\mathbb{R}^{s_{2}\times s_{1}}$ is a $(\epsilon_{2},p_{2},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$ and they are generated independently, we can have
$\frac{\sqrt{1-\epsilon_{s}}}{\sqrt{1+\epsilon_{b}}}>\max(12.8\sqrt{mn{\bf u}+n(n+1){\bf u}}+0.1,0.14)$ (98)
with probability at least $1-p_{f}$. Here, we define $p_{f}=p_{1}+p_{2}-p_{1}p_{2}$, $\epsilon_{s}=\epsilon_{1}+\epsilon_{2}-\epsilon_{1}\epsilon_{2}$ and $\epsilon_{b}=\epsilon_{1}+\epsilon_{2}+\epsilon_{1}\epsilon_{2}$.
With Assumption 1, we provide the primary theorem of rounding error analysis of SSLHC3.
Theorem 4.1 (Rounding error analysis of SSLHC3)
Suppose Assumption 1 is satisfied, if $X\in\mathbb{R}^{m\times n}$, when $[Q_{2ss},R_{4ss}]=SSLHC3(X)$, we can have
$\|Q_{2ss}^{\top}Q_{2ss}-I\|_{F}$$\leq$ $6(mn{\bf u}+n(n+1){\bf u}),$ (99)
$\|Q_{2ss}R_{4ss}-X\|_{F}$$\leq$ $R$ (100)
with probability at least $1-p_{f}$. Here, $R=[1.79(1+h_{1})+4.63\sqrt{1+h_{1}}+1.41]\cdot\frac{h_{2}}{\sqrt{1-\epsilon_{s}}}\cdot n^{2}{\bf u}\|X\|_{2}$, $h_{1}=5\cdot(\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08})^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$ and $h_{2}=\frac{1}{\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.11}{\sqrt{1-\epsilon_{s}}}}$.
For SSLHC3, the following assumptions hold:
$s_{1}\sqrt{n}{\bf u}\leq\frac{1}{64},$ (101)
$s_{2}n^{2}{\bf u}\leq\frac{1}{64},$ (102)
$24.2\cdot\frac{\sqrt{1+\epsilon_{b}}}{\sqrt{1-\epsilon_{s}}}\cdot\sqrt{n}{\bf u}\cdot(s_{1}+s_{2}n\sqrt{n})\kappa_{2}(L)$$\leq$ $1,$ (103)
$22.6n^{2}\sqrt{n}{\bf u}\cdot\frac{\sqrt{1+\epsilon_{b}}}{\sqrt{1-\epsilon_{s}}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\kappa_{2}(L)\kappa_{2}(U)$$\leq$ 1.1 1.1 . (104)
4.2 Theoretical results of SLHC2 and general settings
To analyze SLHC with the single-sketching, we focus on $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding. We provide the following assumption.
Assumption 2 (The assumption for the single-sketching)
If $\Omega\in\mathbb{R}^{s\times m}$ is a $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, we can have
$\frac{\sqrt{1-\epsilon}}{\sqrt{1+\epsilon}}>12.8\sqrt{mn{\bf u}+n(n+1){\bf u}}+0.1$ (105)
with probability at least $1-p$.
With Assumption 2, we provide the primary theorem of rounding error analysis of SLHC2.
Theorem 4.2 (Rounding error analysis of SLHC2)
Suppose Assumption 2 is satisfied, if $X\in\mathbb{R}^{m\times n}$, when $[Q_{1ss},R_{2ss}]=SLHC2(X)$, we can have
$\|Q_{1ss}^{\top}Q_{1ss}-I\|_{F}$$\leq$ $h_{3},$ (106)
$\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $T$ (107)
with probability at least $1-p_{f}$. Here, $T=(2.93\sqrt{1+h_{3}}+1.41)\cdot\frac{h_{4}}{\sqrt{1-\epsilon}}\cdot n^{2}{\bf u}\|X\|_{2}$, $h_{3}=5\cdot(\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon}{1+\epsilon}}-0.08})^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$ and $h_{4}=\frac{1}{\frac{4}{5\sqrt{1+\epsilon}}-\frac{0.11}{\sqrt{1-\epsilon}}}$.
For SLHC2, the following assumptions hold:
$sn^{2}{\bf u}\leq\frac{1}{64},$ (108)
$24.2\cdot\frac{\sqrt{1+\epsilon}}{\sqrt{1-\epsilon}}\cdot\sqrt{n}{\bf u}\cdot(m+sn\sqrt{n})\kappa_{2}(L)$$\leq$ $1,$ (109)
$22.6n^{2}\sqrt{n}{\bf u}\cdot\frac{\sqrt{1+\epsilon}}{\sqrt{1-\epsilon}}\cdot(1+1.08sn\sqrt{n}{\bf u})\kappa_{2}(L)\kappa_{2}(U)$$\leq$ 1.1 1.1 . (110)
4.3 Preparations for proving Theorem 4.1
Regarding Theorem 4.1 and Theorem 4.2, we prove Theorem 4.1 first. The proof of Theorem 4.2 can be received directly from the proof of Theorem 4.1. To prove Theorem 4.1, we provide the following lemmas first.
Lemma 17 (Estimation of $\|E_{Lss}\|_{2}$)
For $\|E_{Lss}\|_{2}$, we can have
$\|E_{Lss}\|_{2}\leq 1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$ (111)
with probability at least $1-p_{f}$.
Proof
To estimate $\|E_{Lss}\|_{2}$, we divide (86) into two steps.
$L_{t}=\Omega_{1}L+E_{L_{1}},$ (112)
$L_{ss}=\Omega_{2}L_{t}+E_{L_{2}}.$ (113)
When we make comparison between (112), (113) and (86), we can have
$E_{Lss}=\Omega_{2}E_{L_{1}}+E_{L_{2}}.$ (114)
Based on Lemma 2 and Lemma 5, we can have
$\|E_{L_{1}}\|_{2}$$\leq$ $\|E_{L_{1}}\|_{F}$ (115)
$\leq$ $\gamma_{m}\cdot\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{1}$. Therefore, with Lemma 2 and (115), we can have
$\|E_{L_{2}}\|_{2}$$\leq$ $\|E_{L_{2}}\|_{F}$ (116)
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{1+\epsilon_{2}}\cdot\|\Omega_{1}L+E_{L_{1}}\|_{F}$
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{1+\epsilon_{2}}\cdot(\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2}+\gamma_{m}\cdot\sqrt{1+\epsilon_{1}}\cdot\sqrt{n}\|L\|_{2})$
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{(1+\epsilon_{1})(1+\epsilon_{2})}\cdot(1+\gamma_{m})\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. With (115), we can have
$\|\Omega_{2}E_{L_{1}}\|_{2}$$\leq$ $\|\Omega_{2}E_{L_{1}}\|_{F}$ (117)
$\leq$ $\gamma_{s_{1}}\cdot\sqrt{1+\epsilon_{2}}\cdot\|E_{L_{1}}\|_{F}$
$\leq$ $\gamma_{s_{1}}\cdot\gamma_{m}\cdot\sqrt{(1+\epsilon_{1})(1+\epsilon_{2})}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. Under the definition in Lemma 7, we put (116) and (117) into (114) and we can have
$\|E_{Lss}\|_{2}$$\leq$ $\|E_{L_{2}}\|_{2}+\|\Omega_{2}E_{L_{1}}\|_{2}$
$\leq$ $\gamma_{s_{1}}\cdot(1+2\gamma_{m})\cdot\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}$
$\leq$ $1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. (111) is proved.
Lemma 18 (Estimation of $\|E_{HQss}\|_{2}$)
For $\|E_{HQss}\|_{2}$, we can have
$\|E_{HQss}\|_{2}\leq 1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$ (118)
with probability at least $1-p_{f}$.
Proof
For $\|E_{HQss}\|_{2}$, with Lemma 3, based on (87), we can have
$\|E_{HQss_{j}}\|_{2}\leq\gamma_{s_{2}n}\cdot\|L_{ss}\|_{2}.$ (119)
Here, $E_{HQss_{j}}$ denotes the $j$-th column of $E_{HQss}$, $j=1,2,\cdots,n$. Based on Lemma 2 and (119), we can have
$\|E_{HQss}\|_{2}$$\leq$ $\|E_{HQss}\|_{F}$ (120)
$\leq$ $\gamma_{s_{2}n}\|L_{ss}\|_{F}$
$\leq$ $1.02s_{2}n\sqrt{n}{\bf u}\|L_{ss}\|_{2}.$
Regarding $\|L_{ss}\|_{2}$, With Lemma 8, (86), (101) and (111), we can get
$\|L_{ss}\|_{2}$$\leq$ $\|\Omega_{2}\Omega_{1}L\|_{2}+\|E_{Lss}\|_{2}$ (121)
$\leq$ $\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}+1.1s_{1}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}$
$\leq$ $1.02\sqrt{1+\epsilon_{b}}\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. We put (121) into (120) and we can get
$\|E_{HQss}\|_{2}$$\leq$ $1.02s_{2}n\sqrt{n}{\bf u}\|L_{ss}\|_{2}$
$\leq$ $1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$
with probability at least $1-p_{f}$. (118) is proved.
Lemma 19 (Estimation of $\|G_{ss}\|_{2}$)
For $\|G_{ss}\|_{2}$, we can have
$\|G_{ss}\|_{2}\leq 1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}$ (122)
with probability at least $1-p_{f}$.
Proof
Regarding $\|G_{ss}\|_{2}$, when $W_{ss}$ is orthogonal, with (87), we can get
$\|G_{ss}\|_{2}$ $=$ $\|W_{ss}G_{ss}\|_{2}$ (123)
$\leq$ $\|L_{ss}\|_{2}+\|E_{HQss}\|_{2}.$
We put (118) and (121) into (123) and we can get
$\|G_{ss}\|_{2}$$\leq$ $\|L_{ss}\|_{2}+\|E_{HQss}\|_{2}$
$\leq$ $1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}$
with probability at least $1-p_{f}$. (122) is proved.
Lemma 20 (Estimation of $\|E_{Rss}\|_{2}$)
For $\|E_{Rss}\|_{2}$, we can have
$\|E_{Rss}\|_{2}\leq 1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$ (124)
with probability at least $1-p_{f}$.
Proof
To estimate $\|E_{Rss}\|_{2}$, based on Lemma 2 and (88), we can get
$\|E_{Rss}\|_{2}\leq 1.1n^{2}{\bf u}\|G_{ss}\|_{2}\|U\|_{2}.$ (125)
We put (122) into (125) and we can get
$\|E_{Rss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|G_{ss}\|_{2}\|U\|_{2}$
$\leq$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. (124) is proved.
Lemma 21 (Estimation of $\|R_{ss}^{-1}\|_{2}$)
For $\|R_{ss}^{-1}\|_{2}$, we can have
$\|R_{ss}^{-1}\|_{2}\leq\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}$ (126)
with probability at least $1-p_{f}$.
Proof
Regarding $\|R_{ss}^{-1}\|_{2}$, with (88), we can have
$\|R_{ss}\|_{2}$ $\leq\|G_{ss}\|_{2}\|U\|_{2}+\|E_{Rss}\|_{2}.$ (127)
We put (122) and (124) into (127) and with (102), we can have
$\|R_{ss}\|_{2}$$\leq$ $\|G_{ss}\|_{2}\|U\|_{2}+\|E_{Rss}\|_{2}$ (128)
$\leq$ $1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}\|U\|_{2}$
$+$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
$\leq$ $1.04\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. With Lemma 4 and (88), we can have
$\sigma_{min}(R_{ss})$ $=$ $\sigma_{min}(G_{ss}U-E_{Rss})$ (129)
$\geq$ $\sigma_{min}(G_{ss})\sigma_{min}(U)-\|E_{Rss}\|_{2}.$
For $\sigma_{min}(G_{ss})$, when $W_{ss}$ is orthogonal, based on Lemma 4 and (87), we can have
$\sigma_{min}(G_{ss})$ $=$ $\sigma_{min}(W_{ss}G_{ss})$ (130)
$\geq$ $\sigma_{min}(L_{ss})-\|E_{HQss}\|_{2}.$
For $\sigma_{min}(L_{ss})$, based on Lemma 4 and (86), we can have
$\sigma_{min}(L_{ss})\geq\sigma_{min}(\Omega_{2}\Omega_{1}L)-\|E_{Lss}\|_{2}.$ (131)
Based on (14) and (86), we can have
$\sigma_{min}(\Omega_{2}\Omega_{1}L)\geq\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)$ (132)
with probability at least $1-p_{f}$. We put (111) and (132) into (131) and we can have
$\sigma_{min}(L_{ss})\geq\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$ (133)
with probability at least $1-p_{f}$. We put (133) into (130) and based on (103), we can have
$\sigma_{min}(G_{ss})$$\geq$ $\sigma_{min}(L_{ss})-\|E_{HQss}\|_{2}$ (134)
$\leq$ $\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)-1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$
-- $1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$
$\geq$ $0.95\cdot\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)$
with probability at least $1-p_{f}$. We put (124) and (134) into (129) and based on (104), we can have
$\sigma_{min}(R_{ss})$$\geq$ $\sigma_{min}(G_{ss})\sigma_{min}(U)-\|E_{Rss}\|_{2}$ (135)
$\geq$ $0.95\cdot\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)\sigma_{min}(U)$
-- $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
$\geq$ $0.95\cdot\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)\sigma_{min}(U)-0.05\cdot\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)\sigma_{min}(U)$
$\geq$ $0.9\cdot\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(L)\sigma_{min}(U)$
with probability at least $1-p_{f}$. With (135), we can estimate $\|R_{ss}^{-1}\|_{2}$ as
$\|R_{ss}^{-1}\|_{2}$ $=$ $\frac{1}{\sigma_{min}(R_{ss})}$
$\leq$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}$
with probability at least $1-p_{f}$. (126) is proved.
Lemma 22 (Estimation of $\|UR_{ss}^{-1}\|_{2}$)
For $\|UR_{ss}^{-1}\|_{2}$, we can have
$\|UR_{ss}^{-1}\|_{2}\leq\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}$ (136)
with probability at least $1-p_{f}$.
Proof
To estimate $\|UR_{ss}^{-1}\|_{2}$, with (88), we can have
$\|UR_{ss}^{-1}\|_{2}\leq\|G^{-1}\|_{2}(1+\|E_{R_{ss}}\|_{2}\|R^{-1}\|_{2}).$ (137)
Regarding $\|G_{ss}^{-1}\|_{2}$, with (134), we can have
$\|G^{-1}\|_{2}$ $=$ $\frac{1}{\sigma_{min}(G_{ss})}$ (138)
$\leq$ $\frac{1.06}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}$
with probability at least $1-p_{f}$. Therefore, we put (124), (126) and (138) into (137) and based on (104), we can have
$\|UR_{ss}^{-1}\|_{2}$$\leq$ $\|G_{ss}^{-1}\|_{2}(1+\|E_{Rss}\|_{2}\|R^{-1}\|_{2})$
$\leq$ $\frac{1.06}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}\cdot(1+(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\cdot\frac{1.25}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)})$
$\leq$ $\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}$
with probability at least $1-p_{f}$. (136) is proved.
Lemma 23 (Estimation of $\|E_{QRss}\|_{2}$)
For $\|E_{QRss}\|_{2}$, we can have
$\|E_{QRss}\|_{2}\leq 1.15n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$ (139)
with probability at least $1-p_{f}$.
Proof
To estimate $\|E_{QRss}\|_{2}$, Based on (89), with Lemma 2, we can have
$\|E_{QRss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|Q_{ss}\|_{2}\|R_{ss}\|_{2}.$ (140)
Regarding $\|R_{ss}\|_{2}$, with (88), we can have
$\|R_{ss}\|_{2}\leq\|G_{ss}\|_{2}\|U\|_{2}+\|E_{R_{ss}}\|_{2}.$ (141)
We put (122) and (124) into (141) and based on (102), we can have
$\|R_{ss}\|_{2}$$\leq$ $\|G_{ss}\|_{2}\|U\|_{2}+\|E_{R_{ss}}\|_{2}$ (142)
$\leq$ $1.02\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\sqrt{n}\|L\|_{2}\cdot\|U\|_{2}$
$+$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
$\leq$ $1.04\sqrt{n}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$
with probability at least $1-p_{f}$. We put (142) into (140) and we can have
$\|E_{QRss}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|Q_{ss}\|_{2}\|R_{ss}\|_{2}$
$\leq$ $1.15n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$
with probability at least $1-p_{f}$. (139) is proved.
4.4 Proof of Theorem Theorem 4.1
With the theoretical results in the previous part, we can prove Theorem 4.1 in this part. Some of the ideas are similar to that of Householder .
Proof
The proof of Theorem 4.1 includes two parts, orthogonality and residual.
4.4.1 Orthogonality
In the beginning, we proof the orthogonality of SSLHC3. With (85), we can have
$\Omega_{2}\Omega_{1}PXR_{ss}^{-1}=\Omega_{2}\Omega_{1}LUR_{ss}^{-1}+E_{LU}R_{ss}^{-1}.$ (143)
Based on (86)-(88), we can have
$\Omega_{2}\Omega_{1}PXR_{ss}^{-1}$ $=$ $\Omega_{2}\Omega_{1}LUR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$ (144)
$=$ $(L_{ss}-E_{L_{ss}})UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $(W_{ss}G_{ss}+E_{HQ_{ss}}-E_{L_{ss}})UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}G_{ss}UR_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}(R_{ss}+E_{Rss})R_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}$
$=$ $W_{ss}+W_{ss}E_{Rss}R_{ss}^{-1}+E_{HQss}UR_{ss}^{-1}-E_{Lss}UR_{ss}^{-1}+E_{LU}R_{ss}^{-1}.$
Based on (144), when $W_{ss}$ is orthogonal, we can have
$\|\Omega_{2}\Omega_{1}PXR_{ss}^{-1}-W_{ss}\|_{2}$$\leq$ $\|E_{Rss}\|_{2}\|R_{ss}^{-1}\|_{2}+\|E_{HQss}\|_{2}\|UR_{ss}^{-1}\|_{2}$ (145)
$+$ $\|E_{Rss}\|_{2}\|UR_{ss}^{-1}\|_{2}+\|E_{LU}\|_{2}\|R_{ss}^{-1}\|_{2}.$
Therefore, we put (56), (118), (124), (126) and (136) into (145) and based on (103) and (104), we can have
$\|\Omega_{2}\Omega_{1}PXR_{ss}^{-1}-W_{ss}\|_{2}$$\leq$ $1.13n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}$ (146)
$\cdot$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}+1.1s_{2}n^{2}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot\|L\|_{2}$
$\cdot$ $\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}+1.1\cdot\sqrt{1+\epsilon_{b}}\cdot s_{1}{\bf u}\cdot\sqrt{n}\|L\|_{2}$
$\cdot$ $\frac{1.12}{\sigma_{min}(L)}\cdot\sqrt{1-\epsilon_{s}}+1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}$
$\cdot$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}$
$\leq$ $\frac{1.12}{20}+\frac{1.12}{22}+\frac{1.24}{24.6}$
$\leq$ 0.20.2 0.20.2
with probability at least $1-p_{f}$. Since $P$ is a permutation matrix, we can easily have
$\sigma_{i}(\Omega_{2}\Omega_{1}XR_{ss}^{-1})=\sigma_{i}(\Omega_{2}\Omega_{1}PXR_{ss}^{-1})$ (147)
when $i=1,2,\cdots,n$. Therefore, with (146) and (147), we can have
$0.8\leq\sigma_{min}(\Omega_{2}\Omega_{1}XR_{ss}^{-1})\leq\|\Omega_{2}\Omega_{1}XR_{ss}^{-1}\|_{2}\leq 1.2$ (148)
with probability at least $1-p_{f}$. Based on (14) and (15), we can have
$\frac{4}{5\sqrt{1+\epsilon_{b}}}\leq\sigma_{min}(XR_{ss}^{-1})\leq\|XR_{ss}^{-1}\|_{2}\leq\frac{6}{5\sqrt{1-\epsilon_{s}}}$ (149)
with probability at least $1-p_{f}$. With (89), we can have
$Q_{ss}-XR_{ss}^{-1}=E_{QRss}R_{ss}^{-1}.$ (150)
Based on (126) and (139), when (104) is satisfied, we can have
$\|Q_{ss}-XR_{ss}^{-1}\|_{2}$$\leq$ $\|E_{QRss}\|_{2}\|R_{ss}^{-1}\|_{2}$ (151)
$\leq$ $1.15n^{2}\sqrt{n}{\bf u}\cdot\sqrt{1+\epsilon_{b}}\cdot(1+1.08s_{2}n\sqrt{n}{\bf u})\cdot\|L\|_{2}\|U\|_{2}\|Q_{ss}\|_{2}$
$\cdot$ $\frac{1.12}{\sqrt{1-\epsilon_{s}}}\cdot\frac{1}{\sigma_{min}(L)\sigma_{min}(U)}$
$\leq$ $0.06\|Q_{ss}\|_{2}$
with probability at least $1-p_{f}$. Based on Lemma 1 and (89), we can have
$\sigma_{min}(XR_{ss}^{-1})-\|Q_{ss}-XR_{ss}^{-1}\|_{2}\leq\sigma_{min}(Q_{ss})\leq\|Q_{ss}\|_{2}\leq\|XR_{ss}^{-1}\|_{2}+\|Q_{ss}-XR_{ss}^{-1}\|_{2}.$ (152)
Based on (149) and (151), we can get $\|Q_{ss}\|_{2}\leq\frac{6}{5\sqrt{1-\epsilon_{s}}}+0.06\|Q_{ss}\|_{2}$ with probability at least $1-p_{f}$. Therefore, it is easy to have
$\|Q_{ss}\|_{2}\leq\frac{1.28}{\sqrt{1-\epsilon_{s}}}$ (153)
with probability at least $1-p_{f}$. With (153), we can have $\|Q_{ss}-XR_{ss}^{-1}\|_{2}\leq\frac{0.08}{\sqrt{1-\epsilon_{s}}}$ with probability at least $1-p_{f}$. Therefore, with (149), we can have
$\sigma_{min}(Q_{ss})\geq\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.08}{\sqrt{1-\epsilon_{s}}}$ (154)
with probability at least $1-p_{f}$. With (153) and (154), we can have
$\kappa_{2}(Q_{ss})$$\leq$ $\frac{\|Q_{ss}\|_{2}}{\sigma_{min}(Q_{ss})}$ (155)
$\leq$ $\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08}$
with probability at least $1-p_{f}$. Under (98), we can easily get $\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08}\leq\frac{1}{8\sqrt{mn{\bf u}+n(n+1){\bf u}}}$. Then based on the theoretical results of error , we can get
$\|Q_{2ss}^{\top}Q_{2ss}-I\|_{F}\leq 6(mn{\bf u}+n(n+1){\bf u})$
with probability at least $1-p_{f}$. Therefore, (99) is proved.
4.4.2 Residual
In the following, we focus on the residual. Regarding (92) and (93), we can have
$Q_{1ss}R_{2ss}$ $=$ $(Q_{ss}+E_{3ss})R_{1ss}^{-1}\cdot(R_{1ss}R_{ss}+E_{4ss})$ (156)
$=$ $(Q_{ss}+E_{3ss})\cdot(R_{ss}+R_{1ss}^{-1}E_{4ss})$
$=$ $Q_{ss}R_{ss}+E_{3ss}R_{ss}+Q_{1ss}E_{4ss}.$
Therefore, with (156), we can have
$\|Q_{1ss}R_{2ss}-X\|_{F}\leq\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}.$ (157)
With Lemma 2, (92) and (93), we can have
$\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}$ (158)
$\leq$ $1.1n^{2}{\bf u}(\|Q_{ss}\|_{2}\|R_{ss}\|_{2}+2\|Q_{1ss}\|_{2}\|R_{1ss}\|_{2}\|R_{ss}\|_{2}).$
For $\|Q_{1ss}R_{2ss}-X\|_{F}$, we can do it in a similar way with that of LUChol . Regarding $\|Q_{1ss}\|_{2}$, with theoretical results in error and (155), we can have
$\|Q_{1ss}^{\top}Q_{1ss}-I\|_{F}$$\leq$ $\frac{5}{64}\cdot(8\kappa_{2}(Q_{ss})\sqrt{mn{\bf u}+n(n+1){\bf u}})^{2}$ (159)
$\leq$ $5\cdot(\frac{1.28}{0.8\cdot\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}-0.08})^{2}\cdot(mn{\bf u}+n(n+1){\bf u})$
$=$ $h_{1}$
with probability at least $1-p_{f}$. Therefore, we can get
$\|Q_{1ss}\|_{2}\leq\sqrt{1+h_{1}}$ (160)
with probability at least $1-p_{f}$. With corresponding results in error and (153), we can have
$\|R_{1ss}\|_{2}$$\leq$ $\frac{65}{63}\|Q_{ss}\|_{2}$ (161)
$\leq$ $\frac{1.33}{\sqrt{1-\epsilon_{s}}}$
with probability at least $1-p_{f}$. Based on (89) and (140), we can have
$\|Q_{ss}R_{ss}-X\|_{2}$$\geq$ $\|Q_{ss}R_{ss}\|_{2}-\|E_{QRss}\|_{2}$ (162)
$\geq$ $\sigma_{min}(Q_{ss})\|R_{ss}\|_{2}-1.1n^{2}\|Q_{ss}\|_{2}\|R_{ss}\|_{2}.$
Therefore, $\|R_{ss}\|_{2}$ can be bounded as
$\|R_{ss}\|_{2}\leq\frac{\|X\|_{2}}{\sigma_{min}(Q_{ss})-1.1n^{2}{\bf u}\|Q_{ss}\|_{2}}.$ (163)
We put (153) and (154) into (163) and based on (102), we can get
$\|R_{ss}\|_{2}$$\leq$ $\frac{\|X\|_{2}}{\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.08}{\sqrt{1-\epsilon_{s}}}-1.1n^{2}{\bf u}\cdot\frac{1.28}{\sqrt{1-\epsilon_{s}}}}$ (164)
$\leq$ $\frac{\|X\|_{2}}{\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.11}{\sqrt{1-\epsilon_{s}}}}$
$=$ $h_{2}\|X\|_{2}$
with probability at least $1-p_{f}$. Here, (98) guarantees $\frac{4}{5\sqrt{1+\epsilon_{b}}}-\frac{0.11}{\sqrt{1-\epsilon_{s}}}>0$. Therefore, We put (153), (160), (161) and (164) into (157) and we can have
$\|Q_{1ss}R_{2ss}-X\|_{F}$$\leq$ $\|Q_{ss}R_{ss}-X\|_{F}+\|E_{3ss}\|_{F}\|R_{ss}\|_{2}+\|E_{4ss}\|_{F}\|Q_{1ss}\|_{2}$ (165)
$\leq$ $(2.93\sqrt{1+h_{1}}+1.41)\cdot\frac{h_{2}}{\sqrt{1-\epsilon_{s}}}\cdot n^{2}{\bf u}\|X\|_{2}$
with probability at least $1-p_{f}$.
Regarding $\|Q_{2}^{\top}R_{4}-X\|_{F}$, with (96) and (97), we can have
$Q_{2ss}R_{4ss}$$\leq$ $(Q_{1ss}+E_{7ss})R_{3ss}^{-1}(R_{3ss}R_{2ss}+E_{8ss})$ (166)
$\leq$ $(Q_{1ss}+E_{7ss})(R_{2ss}+R_{3ss}^{-1}E_{8ss})$
$\leq$ $Q_{1ss}R_{2ss}+E_{7ss}R_{4ss}+Q_{2ss}E_{8ss}.$
Therefore, we can have
$\|Q_{2ss}R_{4ss}-X\|_{F}\leq\|Q_{1ss}R_{2ss}-X\|_{F}+\|E_{7ss}\|_{F}\|R_{4ss}\|_{2}+\|Q_{2ss}\|_{F}\|E_{8ss}\|_{F}.$ (167)
According to the existing results in error , we can have
$\|Q_{2ss}\|_{2}\leq 1.1$ (168)
with probability at least $1-p_{f}$. Moreover, based on (97), we can have
$\|R_{4ss}\|_{2}\leq\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}+\|E_{8ss}\|_{2}.$ (169)
Similar to (169), regarding $\|R_{2ss}\|_{2}$, with (93), we can have
$\|R_{2ss}\|_{2}\leq\|R_{1ss}\|_{2}\|R_{ss}\|_{2}+\|E_{4ss}\|_{2}.$ (170)
With Lemma 2, (102), (161) and (164), we can estimate $\|E_{4ss}\|_{F}$ as
$\|E_{4ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|R_{1ss}\|_{2}\|R_{ss}\|_{2}$ (171)
$\leq$ $\frac{1.1}{64}\cdot\frac{1.33}{\sqrt{1-\epsilon_{s}}}\cdot h_{2}\|X\|_{2}$
with probability at least $1-p_{f}$. Therefore, we put (161), (164) and (171) into (170) and we can get
$\|R_{2ss}\|_{2}$$\leq$ $\|R_{1ss}\|_{2}\|R_{ss}\|_{2}+\|E_{4ss}\|_{2}$ (172)
$\leq$ $\frac{1.36}{\sqrt{1-\epsilon_{s}}}\cdot h_{2}\|X\|_{2}$
with probability at least $1-p_{f}$. Similar to (161) and (160), we can bound $\|R_{3ss}\|_{2}$ as
$\|R_{3ss}\|$$\leq$ $\frac{65}{63}\|Q_{1ss}\|_{2}$ (173)
$\leq$ $\frac{65}{63}\cdot\sqrt{1+h_{1}}$
with probability at least $1-p_{f}$. Therefore, with Lemma 2, (102), (172) and (173), we can estimate $\|E_{8ss}\|_{F}$ as
$\|E_{8ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}$ (174)
$\leq$ $1.1n^{2}{\bf u}\cdot\frac{65}{63}\cdot\frac{1.36}{\sqrt{1-\epsilon_{s}}}\cdot\sqrt{1+h_{1}}\cdot h_{2}\|X\|_{2}$
with probability at least $1-p_{f}$. With Lemma 2, (168) and (173), we can bound $\|E_{7ss}\|_{F}$ as
$\|E_{7ss}\|_{F}$$\leq$ $1.1n^{2}{\bf u}\|Q_{2ss}\|_{2}\|R_{3ss}\|_{2}$ (175)
$\leq$ $1.25n^{2}{\bf u}\cdot\sqrt{1+h_{1}}$
with probability at least $1-p_{f}$. Moreover, with (102), we put (172)-(174) into (169) and we can estimate $\|R_{4ss}\|_{2}$ as
$\|R_{4ss}\|_{2}$$\leq$ $\|R_{3ss}\|_{2}\|R_{2ss}\|_{2}+\|E_{8ss}\|_{2}$ (176)
$\leq$ $\frac{1.43}{\sqrt{1-\epsilon_{s}}}\cdot\sqrt{1+h_{1}}\cdot h_{2}\|X\|_{2}$
with probability at least $1-p_{f}$. In all, we put (165), (168), (174)-(176) into (167) and we can get
$\|Q_{2ss}R_{4ss}-X\|_{F}$$\leq$ $\|Q_{1ss}R_{2ss}-X\|_{F}+\|E_{7ss}\|_{F}\|R_{4ss}\|_{2}+\|Q_{2ss}\|_{2}\|E_{8ss}\|_{F}$
$\leq$ $[1.79(1+h_{1})+4.63\sqrt{1+h_{1}}+1.41]\cdot\frac{h_{2}}{\sqrt{1-\epsilon_{s}}}\cdot n^{2}{\bf u}\|X\|_{2}$
with probability at least $1-p_{f}$. (100) is proved. In all, Theorem 4.1 is proved.
4.5 Proof of Theorem 4.2
Proof
In the proof of Theorem 4.2, we focus directly on (159) and (165). When we consider the single-sketching case with $\Omega=\Omega_{2}$, we can set $\Omega_{1}\in\mathbb{R}^{m\times m}$ to be the identity matrix $I_{m}$, resulting in $s_{1}=m$. Consequently, if $\Omega_{1}\in\mathbb{R}^{s_{1}\times m}$ is a $(\epsilon_{1},p_{1},n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{m}$, we can take $\epsilon_{1}=0$ and $p_{1}=0$. Next, when $\Omega\in\mathbb{R}^{s\times m}$ is a $(\epsilon,p,n)$ oblivious $l_{2}$-subspace embedding in $\mathbb{R}^{s_{1}}$, we let $\epsilon=\epsilon_{b}=\epsilon_{s}$, $p=p_{f}$ and $s=s_{2}$ in (159) and (164). With these settings, we can derive (106) and (107). Thus, Theorem 4.2 is proved.
5 Numerical experiments
In this section, we conduct numerical experiments using MATLAB R2022A on our laptop. Our focus is on three key aspects of the algorithms: their applicability and numerical stability, CPU times, and robustness. We use various examples to test whether our new algorithms can handle a broader range of matrices compared to the original LU-CholeskyQR2. Additionally, we evaluate the numerical stability of these algorithms by examining their orthogonality and residuals. For the CPU time analysis, we compare our new algorithms against the original LU-CholeskyQR2 and CholeskyQR2. We perform numerical experiments with different sizes of input matrices and varying dimensions of sketching. Lastly, the robustness tests are specifically designed for the randomized algorithms, where we assess whether our new randomized algorithms are sufficiently robust for real-world applications.
5.1 The settings and background of our experiments
We show the specification of our computer below in Table 1.
Table 1: The specification of our computer
Item Specification
System Windows 11 family(10.0, Version 22000)
BIOS GBCN17WW
CPU Intel(R) Core(TM) i5-10500H CPU @ 2.50GHz -2.5 GHz
Number of CPUs / node 12
Memory size / node 8 GB
Direct Version DirectX 12
In this section, we conduct numerical experiments using a matrix derived from lower-triangular matrices. Such matrices frequently arise in the fields of numerical PDEs, financial mathematics, and the solution of linear systems (see Practical , Theory , Geometric , Elements , and their additional references). We construct the input matrix $X$ by stacking several small square matrices from top to bottom. Specifically, we utilize a small lower-triangular matrix $X_{1}\in\mathbb{R}^{n\times n}$, defined as follows:
$x_{1ii}$ $=$ $100,i=1,2,\cdots,n,$
$x_{1ij}$ $=$ $a,i=1,2,\cdots,n,j<i,$
$x_{1ij}$ $=$ $0,i=1,2,\cdots,n,j>i,.$
Here, $a$ is a negative constant. $X_{1}$ is a matrix which occurs in some real problems in the industry. We form $X\in\mathbb{R}^{m\times n}$ based on $b$ small $X_{1}$, that is, $X=[X_{1}^{\top},X_{1}^{\top},\cdots,X_{1}^{\top}]^{\top}$. $b$ is a positive integer and $b=\frac{m}{n}$. In the following group of experiments, we very $a$, $b$ and $n$ to do numerical experiments to show the properties of our algorithms. If $X_{ss}=\Omega_{2}\Omega_{1}X$, $X\in\mathbb{R}^{m\times n}$, we let $\epsilon_{1}=0.5$, $p_{1}=0.6$, $\epsilon_{2}=0.5$ and $p_{2}=0.4$ in Definition 2. We take different $m$, $n$, $s_{1}$ and $s_{2}$ in the following experiments. Therefore, $\epsilon_{b}=1.25$, $\epsilon_{s}=0.75$. We can easily find that Assumption 1 is satisfied if $mn$ do not exceed the level of $10^{12}$. This is enough for the problems in the engineering. Regarding the single-sketching, we take the Gaussian sketch. Here, we have $\epsilon=0.5$ and $p=0.6$. When $mn$ do not exceed the level of $10^{12}$, our settings can also meet the requirement in Assumption 2. Readers are encouraged to conduct experiments with other types of sketches as well.
5.2 Tests of the applicability and the numerical stability of the algorithms
In this group of experiments, we focus on the applicability and numerical stability of our algorithms, comparing them with the original LU-CholeskyQR2. First, we examine the potential influence of $\kappa_{2}(X)$. We conduct two groups of numerical experiments with different values of $m$: $m=20000$ and $m=30000$. We fix $n=50$ and vary $a$ from $-70$, $-80$, $-90$ to $-100$ to alter $\kappa_{2}(X)$. For the sizes of $\Omega_{1}$ and $\Omega_{2}$, we set $s_{1}=17000$ and $s_{2}=50$ for SLHC2, and $s=50$ for SSLHC3, based on (9) and (10). Our numerical experiments are summarized in Tables 24, and Table 6. Next, we shift our focus to $n$. We fix $a=70$ and vary $n$ from $5$, $10$, $20$ to $50$ while conducting numerical experiments. Here, we set $s_{1}=\frac{n^{2}+n}{0.5\cdot 0.6^{2}}=\frac{n^{2}+n}{0.15}$ and $s_{2}=50$ for SSLHC3, while $s=50$ for SLHC2. We perform two groups of numerical experiments with $m=20000$ and $m=30000$. The results of these experiments are listed in Table 6 and Table 7.
According to Tables 27, we find that our LHC2, SLHC2, and SSLHC3 demonstrate significant advantages over the original LU-CholeskyQR2 in terms of applicability when $X$ is ill-conditioned. This is a key innovative point of our work. Among the three new algorithms, our deterministic LHC2 exhibits good numerical stability in both orthogonality and residuals. However, the orthogonality of SLHC2 is less stable and relatively worse than that of LHC2 and SSLHC3. This instability arises because SLHC2 lacks the CholeskyQR2 structure that underpins SLHC. This can also be observed in the comparisons between (25), (99), and (106). In the randomized algorithms, the orthogonality and residuals of SSLHC3 are nearly on par with those of the deterministic LU-CholeskyQR2 and LHC2. This demonstrates the effectiveness of multi-sketching in ensuring numerical stability. Our SSLHC3 effectively balances numerical stability and applicability.
Table 2: Comparison of orthogonality with $\kappa_{2}(X)$ varying when $m=20000$ and $n=50$
$\kappa_{2}(X)$ $122.65e+$ 12 $135.10e+$ 13 $148.29e+$ 14 $161.13e+$ 16
LU-CholeskyQR2 $156.52e-$ 15 -- -- --
LHC2 $159.19e-$ 15 $155.52e-$ 15 $158.79e-$ 15 $159.67e-$ 15
SLHC2 $145.50e-$ 14 $134.19e-$ 13 $147.76e-$ 14 $145.94e-$ 14
SSLHC3 $159.05e-$ 15 $157.10e-$ 15 $158.54e-$ 15 $159.37e-$ 15
Table 3: Comparison of residual with $\kappa_{2}(X)$ varying when $m=20000$ and $n=50$
$\kappa_{2}(X)$ $122.65e+$ 12 $135.10e+$ 13 $148.29e+$ 14 $161.13e+$ 16
LU-CholeskyQR2 $111.15e-$ 11 -- -- --
LHC2 $111.34e-$ 11 $111.64e-$ 11 $111.87e-$ 11 $111.63e-$ 11
SLHC2 $111.68e-$ 11 $112.27e-$ 11 $112.26e-$ 11 $112.14e-$ 11
SSLHC3 $112.29e-$ 11 $112.46e-$ 11 $112.78e-$ 11 $112.79e-$ 11
Table 4: Comparison of orthogonality with $\kappa_{2}(X)$ varying when $m=30000$ and $n=50$
$\kappa_{2}(X)$ $122.65e+$ 12 $135.10e+$ 13 $148.29e+$ 14 $161.12e+$ 16
LU-CholeskyQR2 $154.82e-$ 15 -- -- $159.07e-$ 15
LHC2 $155.76e-$ 15 $158.59e-$ 15 $155.92e-$ 15 $141.02e-$ 14
SLHC2 $131.08e-$ 13 $146.93e-$ 14 $131.40e-$ 13 $132.24e-$ 13
SSLHC3 $155.71e-$ 15 $157.09e-$ 15 $159.90e-$ 15 $159.81e-$ 15
Table 5: Comparison of residual with $\kappa_{2}(X)$ varying when $m=30000$ and $n=50$
$\kappa_{2}(X)$ $122.65e+$ 12 $135.10e+$ 13 $148.29e+$ 14 $161.13e+$ 16
LU-CholeskyQR2 $111.50e-$ 11 -- -- $111.74e-$ 11
LHC2 $111.81e-$ 11 $111.75e-$ 11 $111.81e-$ 11 $112.22e-$ 11
SLHC2 $112.38e-$ 11 $112.15e-$ 11 $113.30e-$ 11 $112.66e-$ 11
SSLHC3 $112.98e-$ 11 $112.54e-$ 11 $113.17e-$ 11 $113.18e-$ 11
Table 6: Comparison of residual with different $n$ varying when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
LU-CholeskyQR2 $153.85e-$ 15 $156.02e-$ 15 $152.38e-$ 15 --
LHC2 $154.36e-$ 15 $159.60e-$ 15 $156.73e-$ 15 $159.67e-$ 15
SLHC2 $154.58e-$ 15 $141.01e-$ 14 $141.07e-$ 14 $145.94e-$ 14
SSLHC3 $153.05e-$ 15 $154.67e-$ 15 $154.86e-$ 15 $157.64e-$ 15
Table 7: Comparison of residual with different $n$ varying when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
LU-CholeskyQR2 $121.75e-$ 12 $124.46e-$ 12 $127.00e-$ 12 --
LHC2 $122.47e-$ 12 $124.91e-$ 12 $127.63e-$ 12 $111.63e-$ 11
SLHC2 $122.43e-$ 12 $124.18e-$ 12 $125.70e-$ 12 $112.14e-$ 11
SSLHC3 $123.58e-$ 12 $124.65e-$ 12 $124.30e-$ 12 $111.88e-$ 11
5.3 Tests of CPU times of the algorithms
In this section, we evaluate the CPU times of the algorithms presented in this work. Several factors influence the CPU times, so we compare results based on $n$ and $\kappa_{2}(X)$. Additionally, we conduct a group of experiments to test the influence of the size of the Gaussian sketch between SLHC2 and SSLHC3. We examine cases where $m=20000$ and $m=30000$ in each group of experiments. For varying $n$, we fix $a=-70$. For SLHC2, we set $s=50$. For SSLHC3, we fix $s_{2}=50$ and set $s_{1}=\frac{n^{2}+n}{0.15}$. We vary $n$ from $5$, $10$, $20$ to $50$. The numerical experiments testing the CPU times of our three algorithms are listed in Table 8 and Table 9. Regarding the size of the Gaussian sketch, we fix $n=20$ and $a=-70$. For SLHC2, we vary $s$, while for SSLHC3, we fix $s_{1}=2800$ and vary $s_{2}$. We change $s$ from $50$, $100$, $200$ to $400$ and conduct numerical experiments. The results are presented in Table 10 and Table 11. Next, we analyze the influence of $\kappa_{2}(X)$ on CPU times, fixing $n=20$. For SLHC2, we set $s=50$, and for SSLHC3, we maintain $s_{1}=2800$ and $s_{2}=50$. We vary $a$ from $-70$, $-80$, $-90$ to $-100$ to explore the effects on $\kappa_{2}(X)$. Table 12 and Table 13 reflect the CPU times of SLHC2 and SSLHC3 under different $\kappa_{2}(X)$ values. Tables 1421 compare the CPU times of our two randomized algorithms with some deterministic CholeskyQR-type algorithms across various cases. To examine the influence of different $n$, we fix $a=-10$. For SLHC2, we set $s=50$, while for SSLHC3, we fix $s_{2}=50$ and calculate $\frac{n^{2}+n}{0.15}$. We vary $n$ from $5$, $10$, $20$ to $50$. When varying $\kappa_{2}(X)$, we fix $n=20$. For SLHC2, we set $s=50$, and for SSLHC3, we set $s_{1}=2800$ and $s_{2}=50$. We vary $a$ from $-10$, $-20$, $-30$ to $-40$. Tables 14-17 include the CPU times of SLHC2, CholeskyQR2, and LU-CholeskyQR2, while Tables 1821 present the CPU times of SSLHC3 and Shifted CholeskyQR3.
Table 8 and Table 9 demonstrate that our randomized algorithms are significantly more efficient than LHC2 across different values of $m$ and $n$. This represents the primary advantage of randomized algorithms over deterministic ones. Based on the results in Tables 813, we find that our SSLHC3 with multi-sketching is also faster than SLHC2 with single-sketching in many comparison cases, highlighting the unique value of the multi-sketching technique compared to the original single-sketching approach. When compared to other deterministic CholeskyQR-type algorithms, our SSLHC3 operates at a speed comparable to LU-CholeskyQR2, while both CholeskyQR2 and Shifted CholeskyQR3 are faster. However, our randomized algorithms offer better applicability than these two deterministic algorithms. Overall, we can conclude that our SLHC2 and SSLHC3 are efficient. The multi-sketching technique proves to be a superior method for accelerating algorithms.
Table 8: Comparison of CPU time(s) between different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
LHC2 $8.300$ $10.778$ $19.612$ $45.667$
SLHC2 $0.030$ $0.031$ $0.039$ $0.087$
SSLHC3 $0.022$ $0.027$ $0.037$ $0.084$
Table 9: Comparison of CPU time(s) between different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
LHC2 $130.643$ $264.112$ $520.114$ $1396.200$
SLHC2 $0.048$ $0.052$ $0.061$ $0.163$
SSLHC3 $0.037$ $0.044$ $0.060$ $0.145$
Table 10: Comparison of CPU time(s) with different $s$ for the Gaussian sketch when $m=20000$ and $n=20$
$s$ $50$ $100$ $200$ $400$
SLHC2 $0.040$ $0.052$ $0.088$ $0.117$
SSLHC3 $0.038$ $0.040$ $0.040$ $0.050$
Table 11: Comparison of CPU time(s) with different $s$ for the Gaussian sketch when $m=30000$ and $n=20$
$s$ $50$ $100$ $200$ $400$
SLHC2 $0.061$ $0.080$ $0.102$ $0.197$
SSLHC3 $0.060$ $0.055$ $0.059$ $0.065$
Table 12: Comparison of CPU time(s) with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
$\kappa_{2}(X)$ $051.23e+$ 05 $054.28e+$ 05 $061.38e+$ 06 $064.15e+$ 06
SLHC2 $0.040$ $0.052$ $0.088$ $0.117$
SSLHC3 $0.038$ $0.040$ $0.040$ $0.050$
Table 13: Comparison of CPU time(s) with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
$\kappa_{2}(X)$ $051.23e+$ 05 $054.28e+$ 05 $061.38e+$ 06 $064.15e+$ 06
SLHC2 $0.061$ $0.080$ $0.102$ $0.197$
SSLHC3 $0.060$ $0.055$ $0.059$ $0.065$
Table 14: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
CholeskyQR2 $0.002$ $0.006$ $0.009$ $0.024$
LU-CholeskyQR2 $0.025$ $0.025$ $0.031$ $0.050$
SLHC2 $0.030$ $0.033$ $0.038$ $0.059$
Table 15: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
CholeskyQR2 $0.003$ $0.005$ $0.013$ $0.037$
LU-CholeskyQR2 $0.037$ $0.036$ $0.048$ $0.089$
SLHC2 $0.050$ $0.052$ $0.058$ $0.095$
Table 16: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
$\kappa_{2}(X)$ $4.65$ $38.61$ $274.50$ $031.57e+$ 03
CholeskyQR2 $0.008$ $0.009$ $0.007$ $0.009$
LU-CholeskyQR2 $0.029$ $0.032$ $0.030$ $0.046$
SLHC2 $0.041$ $0.039$ $0.041$ $0.040$
Table 17: Comparison of CPU time(s) between SLHC2 and other two-step algorithms with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
$\kappa_{2}(X)$ $4.65$ $38.61$ $274.50$ $031.57e+$ 03
CholeskyQR2 $0.017$ $0.013$ $0.015$ $0.016$
LU-CholeskyQR2 $0.052$ $0.051$ $0.052$ $0.055$
SLHC2 $0.065$ $0.059$ $0.068$ $0.077$
Table 18: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $n$ when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
Shifted CholeskyQR3 $0.002$ $0.006$ $0.014$ $0.039$
SSLHC3 $0.022$ $0.027$ $0.037$ $0.084$
Table 19: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $n$ when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
$n$ $5$ $10$ $20$ $50$
Shifted CholeskyQR3 $0.004$ $0.012$ $0.024$ $0.066$
SSLHC3 $0.037$ $0.044$ $0.060$ $0.145$
Table 20: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $\kappa_{2}(X)$ when $m=20000$ and $n=20$
$\kappa_{2}(X)$ $4.65$ $38.61$ $274.50$ $031.57e+$ 03
Shifted CholeskyQR3 $0.014$ $0.010$ $0.016$ $0.018$
SSLHC3 $0.039$ $0.036$ $0.038$ $0.039$
Table 21: Comparison of CPU time(s) between SSLHC3 and Shifted CholeskyQR3 with different $\kappa_{2}(X)$ when $m=30000$ and $n=20$
$\kappa_{2}(X)$ $4.65$ $38.61$ $274.50$ $031.57e+$ 03
Shifted CholeskyQR3 $0.024$ $0.023$ $0.026$ $0.057$
SSLHC3 $0.060$ $0.058$ $0.064$ $0.025$
5.4 Tests of the robustness of the randomized algorithms
In this section, we present a series of distinct numerical experiments that, to our knowledge, have not been conducted in other similar works. We will assess the robustness of our two randomized algorithms after performing a certain number of experiments. Specifically, we will examine the influence of $\kappa_{2}(X)$, the size of the sketching matrix, and $n$ on the robustness of the algorithms. We conduct two groups of experiments with $m=20000$ and $m=30000$ in each case. To test the effect of different $\kappa_{2}(X)$, we fix $n=50$. For SLHC2, we set $s=50$. For SSLHC3, we set $s_{1}=17000$ and $s_{2}=50$. We vary $a$ from $-70$, $-80$, $-90$ to $-100$ to explore the impact on $\kappa_{2}(X)$. We record the number of successful outcomes every $30$ trials after conducting several groups and calculate the average. The numerical results are presented in Table 22 and Table 23. Regarding the size of the sketching matrix, we focus on the Gaussian sketch, which corresponds to $s$ in SLHC2 and $s_{2}$ in SSLHC3. We fix $n=50$ and $a=-100$, varying the size of the Gaussian sketch from $50$, $100$, $200$ to $400$. The average numbers of successful results in $30$ trials after several groups of experiments are listed in Table 24 and Table 25. When considering different values of $n$, we fix $a=-100$. For SLHC2, we maintain $s=50$. For SSLHC3, we fix $s_{2}=50$ and set $s_{1}=\frac{n^{2}+n}{0.15}$. We vary $n$ from $5$, $10$, $20$ to $50$. The numerical results are detailed in Table 26 and Table 27.
The numerical results in Tables 2227 demonstrate that both our SLHC2 and SSLHC3 are robust enough for real applications across different variants. Based on all the numerical experiments, SSLHC3 is efficient and accurate, exhibiting excellent applicability and robustness. It represents an ideal CholeskyQR-type algorithm that does not require $\kappa_{2}(X)$.
Table 22: Times of success in $30$ times with different $\kappa_{2}(X)$ when $m=20000$ and $n=50$
$\kappa_{2}(X)$ $122.65e+$ 12 $135.10e+$ 13 $148.29e+$ 14 $161.13e+$ 16
SLHC2 $30$ $30$ $30$ $30$
SSLHC3 $30$ $30$ $30$ $30$
Table 23: Times of success in $30$ times with different $\kappa_{2}(X)$ when $m=30000$ and $n=50$
$\kappa_{2}(X)$ $122.65e+$ 12 $135.10e+$ 13 $148.29e+$ 14 $161.13e+$ 16
SLHC2 $30$ $30$ $30$ $30$
SSLHC3 $30$ $30$ $30$ $30$
Table 24: Times of success in $30$ times with different $s$ of the Gaussian sketch when $m=20000$ and $n=50$
s $50$ $100$ $200$ $400$
SLHC2 $30$ $30$ $30$ $30$
SSLHC3 $30$ $30$ $30$ $30$
Table 25: Times of success in $30$ times with different $s$ of the Gaussian sketch when $m=30000$ and $n=50$
s $50$ $100$ $200$ $400$
SLHC2 $30$ $30$ $30$ $30$
SSLHC3 $30$ $30$ $30$ $30$
Table 26: Times of success in $30$ times with different $n$ of the Gaussian sketch when $m=20000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
n $5$ $10$ $20$ $50$
SLHC2 $30$ $30$ $30$ $30$
SSLHC3 $30$ $30$ $30$ $30$
Table 27: Times of success in $30$ times with different $n$ of the Gaussian sketch when $m=30000$ and $s_{1}=\frac{n^{2}+n}{0.15}$
n $5$ $10$ $20$ $50$
SLHC2 $30$ $30$ $30$ $30$
SSLHC3 $30$ $30$ $30$ $30$
6 Conclusions
In this work, we focus on improving the LU-CholeskyQR2 algorithm. Our new deterministic method, LHC2, employs a different approach to generating the $R$-factor compared to LU-CholeskyQR2, which helps avoid numerical breakdown in certain ill-conditioned cases, thereby enhancing the applicability of the algorithm. We utilize the latest randomized technique, matrix sketching, to accelerate our new algorithms during the HouseholderQR step and propose two randomized algorithms: SLHC2 and SSLHC3. We conduct a detailed rounding error analysis of all three new algorithms. Numerical experiments demonstrate that LHC2, SLHC2, and SSLHC3 all exhibit better and more satisfactory performance in terms of applicability compared to the original LU-CholeskyQR2. Furthermore, SLHC2 and SSLHC3 are significantly more efficient than LHC2, highlighting the advantages of the randomized technique. In particular, SSLHC3, which incorporates multi-sketching, outperforms SLHC2 and effectively balances applicability, accuracy, efficiency, and robustness without requiring $\kappa_{2}(X)$. This makes it one of the best algorithms among CholeskyQR-type algorithms.
7 Acknowledgement
This work is supported by the CAS AMSS-PolyU Joint Laboratory of Applied Mathematics. The research of Haoran Guan is funded by the Hong Kong Research Grants Council under RFS grant RFS2021-5S03 and GRF grant 15302122, as well as by the Hong Kong Polytechnic University grant 4-ZZLS. We would like to acknowledge the support of Professor Zhonghua Qiao from the Hong Kong Polytechnic University. We are also grateful to Mr. Yuan Liang from Beijing Normal University, Zhuhai, for his valuable suggestions and examples related to the numerical experiments.
Conflict of interest
The authors declare that they have no conflict of interest.
Data availability
The authors declare that all data supporting the findings of this study are available within this article.